### v2

# Anderson localization on quantum graphs coded by elements of a subshift of finite type

Abstract.
We study Schrödinger operators on quantum graphs where the number of edges between points is determined by orbits of a ”shift of finite type”. We prove Anderson localization for these systems.
1. Main result
This paper extends findings from [27] regarding the positivity of the Lyapunov exponent in quantum graphs with a variable number of edges. By combining this positivity with large deviation estimates, we establish spectral characteristics and prove localization for these systems.
For a positive integer $\ell>1$, let $\mathcal{A}$$^{\mathbb{Z}}$ be the collection of all infinite sequences $\{\omega_{n}\}_{n\in{\mathbb{Z}}}$ such that $\omega_{n}\in{\mathcal{A}}$, where ${\mathcal{A}}=\{1,\dots,\ell\}$. To make it more complicated, we will exclude certain sequences in $\mathcal{A}$$^{\mathbb{Z}}$ from consideration. Namely, after fixing a certain set ${\mathcal{F}}\subset{\mathcal{A}}\times{\mathcal{A}}$, we consider the collection $\Omega\subset{\mathcal{A}}^{\mathbb{Z}}$ of all infinite sequences $\{\omega_{n}\}_{n\in{\mathbb{Z}}}$ for which
$(\omega_{n},\omega_{n+1})\notin{\mathcal{F}},\qquad\forall n\in{\mathbb{Z}}.$
It is easy to see that $\Omega$ is a compact metric space with respect to the metric $d(\cdot,\cdot)$ defined by
$d(\omega,\omega^{\prime})=e^{-N(\omega,\omega^{\prime})},$
where $N(\omega,\omega^{\prime})$ is the largest nonnegative integer such that $\omega_{n}=\omega_{n}^{\prime}$ for all $|n|<N(\omega,\omega^{\prime})$. Define the mapping $T:\,\Omega\to\Omega$ by
$\bigl(T\omega\bigr)_{n}=\omega_{n+1},\qquad\forall n\in{\mathbb{Z}}.$
Such a mapping $T$ is called a subshift of finite type. If $\mathcal{F}$$\emptyset$, then $T$ is called the full shift.
For each $\omega\in\Omega$, we construct the graph $\Gamma_{\omega}$, displayed below for the case where $\ell=2$, $\mathcal{F}$$\emptyset$, and $\omega=\dots 2,1,1,2,2,2,1,1,2,1,2,1,1,1,\dots$
2211112222221111221122111111
Namely, let $\mathbb{Z}$} be the set of integer numbers. For each $\omega\in\Omega$ and $n\in{\mathbb{Z}}$, we consider $\omega_{n}$ copies of the interval $[n,n+1]$. Denoting these copies by $I_{n,j}$, where $j=1,\dots,\omega_{n}$, we define the graph $\Gamma_{\omega}$ as the union
$\Gamma_{\omega}=\bigcup_{n\in{\mathbb{Z}}}\Bigl(\bigcup_{j=1}^{\omega_{n}}I_{n,j}\Bigr).$ (1.1)
While the interiors of the intervals $I_{n,j}$ are assumed to be disjoint, we will also assume that their endpoints are shared in the sense that $n$ and $n+1$ belong to all intervals $I_{n,j}$. Thus,
$\bigcap_{j=1}^{\omega_{n}}I_{n,j}=\{n\}\cup\{n+1\}.$ (1.2)
There is a natural Lebesgue measure on $\Gamma_{\omega}$ whose restriction to $I_{n,j}$ is the Lebesgue measure on this interval. The main object of our study is the Schrödinger operator $H_{\omega}$ formally defined by
$H_{\omega}u=-u^{\prime\prime}$ (1.3)
on the domain $D(H_{\omega})$ that is contained in the orthogonal sum of Sobolev spaces $W^{2,2}(I_{n,j})$. Besides the condition
$D(H_{\omega})\subset\bigoplus_{n=-\infty}^{\infty}\bigoplus_{j=1}^{\omega_{n}}\quad W^{2,2}(I_{n,j}),$
we assume that all functions $u\in D(H_{\omega})$ are continuous on $\Gamma_{\omega}$, and the restrictions $u_{n,j}$ of $u$ to $I_{n,j}$ satisfy
$\sum_{j=1}^{\omega_{n}}u^{\prime}_{n,j}(n)=\sum_{j=1}^{\omega_{n-1}}u^{\prime}_{n-1,j}(n),\quad\forall n\in{\mathbb{Z}}.$ (1.4)
The last relation is called Kirchhoff’s gluing condition at the point $n$. Note that the operator $H_{\omega}$ is self-adjoint in the space $L^{2}(\Gamma_{\omega})$.
Since $\Omega$ is a metric space, we can talk about the Borel $\sigma$-algebra of subsets of $\Omega$ and consider probability measures on $\Omega$. Let $\mu$ be a $T$-ergodic probability measure on $\Omega$. We will not work with arbitrary measures $\mu$. Instead, we will assume that $\mu$ has a bounded distortion property. To give a formal definition of this property, we first define the cylinder sets
$n;j_{0},\dots,j_{k}]=\{\omega\in\Omega:\,\,\omega_{n+s}=j_{s},\,\,s=0,\dots,k\}$
Definition. The measure $\mu$ is said to be a measure with a bounded distortion property, provided there is a constant $C>1$ such that
$C^{-1}\leqslant\frac{\mu([n;j_{0},\dots j_{k}]\cap[l;i_{0},\dots i_{s}])}{\mu([n;j_{0},\dots j_{k}])\cdot\mu([l;i_{0},\dots i_{s}])}\leqslant C$
for all $l>n+k$ and $n;j_{0},\dots j_{k}]\cap[l;i_{0},\dots i_{s}]\neq\emptyset$.
Our main result is the following theorem.
Theorem 1.1.
Let $\Omega$ be a subshift of finite type and let $\mu$ be a $T$-ergodic measure with the bounded distortion property such that ${\rm supp}\,(\mu)=\Omega$. Suppose $T$ has at least one fixed point and at least one non-fixed point. Then for $\mu$-almost every $\omega$, the operator $H_{\omega}$ has pure point spectrum equal to $0,\infty)$. Furthermore, there is a finite subset ${\mathfrak{X}}\subset[0,2\pi]$, independent of $\omega$, such that for any eigenvalue $E$ of $H_{\omega}$, if $\sqrt{E}-2\pi k\notin{\mathfrak{X}}$ for all $k=0,1,\dots$, then the corresponding eigenfunction decays exponentially as $|x|\to\infty.$
This theorem can be viewed as an analogue of the result by Avila, Damanik, and Zhang [2] in which their framework is changed from standard one-dimensional discrete Schrödinger operators to a different setting of quantum graphs while maintaining similar underlying dynamics. Although we adopt the methodology developed by Avila, Damanik, and Zhang [2], certain elements of our proofs differ to accommodate our specific setting.
2. Positivity of the Lyapunov exponent. Large deviations
The proof of Theorem 1.1 is based on positivity of the Lyapunov exponent established in [27]. Note that the conditions on $\mu$ in [27] were weaker than in Theorem 1.1. We assumed that $\mu$ has a local product structure. Let us now give a formal definition of a measure having this property. We first define the spaces of semi-infinite sequences
[…]
Then using the natural projection $\pi_{\pm}$ from $\Omega$ onto $\Omega_{\pm}$, we define $\mu_{\pm}=(\pi_{\pm})_{*}\mu$ on $\Omega_{\pm}$ to be the pushforward measures of $\mu$. After that, for each $1\leqslant j\leqslant\ell$, we introduce the cylinder sets
$0;j]=\{\omega\in\Omega:\,\,\omega_{0}=j\}\quad\text{and}\quad[0;j]_{\pm}=\{\omega\in\Omega_{\pm}:\,\,\omega_{0}=j\}.$
A local product structure is a relation between the measures $\mu_{j}=\mu\bigl|_{[0;j]}$ and the measures $\mu_{j}^{\pm}=\mu_{\pm}\bigl|_{[0;j]}$. To describe this relation, we need to consider the natural homeomorphisms
$P_{j}:[0;j]\to[0;j]_{-}\times[0;j]_{+}$
defined by
$P_{j}(\omega)=\bigl(\pi_{-}\omega,\pi_{+}\omega\bigr),\qquad\forall\omega\in\Omega.$
[…]
Proposition 2.1.
Let $k\neq\pi j$ for all $j\in{\mathbb{Z}}$. Let $u\in C(\Gamma_{\omega})$ be an absolutely continuous solution of the equation
$u^{\prime\prime}(x)=k^{2}u(x),\qquad\text{for a.e. }\quad x\in\Gamma_{\omega}.$ (2.1)
satisfying Kirchhoff’s condition (1.4) at each $n\in{\mathbb{Z}}$. Then
$\omega_{n}u(n+1)+\omega_{n-1}u(n-1)-(\omega_{n}+\omega_{n-1})\cos(k)u(n)=0.$ (2.2)
We will say that $E=k^{2}$ is a generalized eigenvalue for $H_{\omega}$, if (2.1) has a nontrivial solution obeying
$|u(n)|\leqslant C_{u}(1+|n|),\qquad\forall n\in{\mathbb{Z}}.$ (2.3)
A solution satisfying (2.3) is called a generalized eigenfunction.
It turns out that spectral properties of $H_{\omega}$ are closely related to that of the operator ${\mathcal{H}}_{\omega}$ which is defined on $\ell^{2}({\mathbb{Z}})$ by
$\bigl[{\mathcal{H}}_{\omega}u\bigr](n)=\frac{2\omega_{n}u(n+1)}{\sqrt{(\omega_{n+1}+\omega_{n})(\omega_{n}+\omega_{n-1})}}+\frac{2\omega_{n-1}u(n-1)}{\sqrt{(\omega_{n-1}+\omega_{n-2})(\omega_{n}+\omega_{n-1})}}$ (2.4)
for each $u\in\ell^{2}({\mathbb{Z}})$. Generalized eigenvalues and generalized eigenfunctions for ${\mathcal{H}}_{\omega}$ are introduced in the same way as for $H_{\omega}$.
Proposition 2.2.
Let $k\neq\pi n$ for all $n\in{\mathbb{Z}}$. The point $E=k^{2}>0$ is a generalized eigenvalie of the operator $H_{\omega}$ if and only if $2\cos k$ is a generalized eigenvalie of the operator ${\mathcal{H}}_{\omega}$.
Spectral properties of $H_{\omega}$ are related to the behavior of solutions to the equation (2.2). On the other hand, all solutions to (2.2) can be described in terms of the cocycles $(T,A^{E})$ with $A^{E}:\,\Omega\to{\rm SL}(2,{\mathbb{R}})$ defined by
$A^{E}(\omega)=\sqrt{\frac{\omega_{0}}{\omega_{-1}}}\begin{pmatrix}\frac{\omega_{0}+\omega_{-1}}{\omega_{0}}\cos(k)&-\frac{\omega_{-1}}{\omega_{0}}\\ 1&0\end{pmatrix},\qquad E=k^{2}.$ (2.5)
Namely, $u$ is a solution of (2.2) if and only if
$\begin{pmatrix}u(n)\\ u(n-1)\end{pmatrix}=\sqrt{\frac{\omega_{-1}}{\omega_{n-1}}}A^{E}_{n}(\omega)\cdot\begin{pmatrix}u(0)\\ u(-1)\end{pmatrix},\qquad\forall n\in{\mathbb{Z}},$
where
$A^{E}_{n}(\omega)=\begin{cases}A^{E}(T^{n-1}\omega)\cdots A^{E}(\omega)\quad\text{if}\quad n\geqslant 1;\\ [A^{E}_{-n}(T^{n}\omega)]^{-1}\quad\text{if}\quad n\leqslant-1;\\ {\rm Id}\quad\text{if}\quad n=0.\end{cases}$
The Lyapunov exponent for $A^{E}$ and $\mu$ is defined by
$L(E)=\lim_{n\to\infty}\frac{1}{n}\int\ln(\|A^{E}_{n}(\omega)\|)d\mu(\omega).$
Clearly, $L(E)\geqslant 0$. By Kingman’s subaddive ergodic theorem,
$\frac{1}{n}\ln(\|A^{E}_{n}(\omega)\|)\qquad\text{converges to}\quad L(E)\qquad\text{as}\quad n\to\infty,$
for $\mu$-almost every $\omega\in\Omega$.
The main result of [27] desribes the properties of the set
${\mathfrak{L}}(\mu)=\{E\in[0,4\pi^{2}]:\,\,L(E)=0\}.$
Theorem 2.3.
Let $T:\Omega\to\Omega$ be a subshift of finite type and $\mu$ be a $T$-ergodic probability measure with full support (${\rm supp}\,\mu=\Omega$) and a local product structure. Assume $T$ possesses at least one fixed point and at least one non-fixed point $\omega\in\Omega$. Then the set ${\mathfrak{L}}(\mu)$ is finite. Furthermore, if $E>0$ satisfies $(\sqrt{E}-2\pi n)^{2}\notin{\mathfrak{L}}(\mu)$ for all $n=0,1,\dots$, then $L(E)>0$.
Definition . Let $T:\Omega\to\Omega$ be a subshift of finite type. The local stable set of a point $\omega\in\Omega$ is defined by
$W_{\rm loc}^{s}(\omega)=\{\omega^{\prime}\in\Omega:\,\,\omega^{\prime}_{n}=\omega_{n}\quad\text{for}\quad n\geqslant 0\}$
and the local unstable set of $\omega$ is defined by
$W_{\rm loc}^{u}(\omega)=\{\omega^{\prime}\in\Omega:\,\,\omega^{\prime}_{n}=\omega_{n}\quad\text{for}\quad n\leqslant 0\}.$
For $\omega^{\prime}\in W_{\rm loc}^{s}(\omega)$, define $H_{\omega^{\prime},\omega}^{s,n}$ to be
$H_{\omega,\omega^{\prime}}^{s,n}=\bigl[A^{E}_{n}(\omega^{\prime})\bigr]^{-1}A^{E}_{n}(\omega).$
Since $d(T^{j}\omega^{\prime},T^{j}\omega)\leqslant e^{-j}$ tends to 0 as $j\to\infty$, there is an index $n_{0}$ for which
$H_{\omega,\omega^{\prime}}^{s,n}=H_{\omega,\omega^{\prime}}^{s,n_{0}}\qquad\text{for }\quad n\geqslant n_{0}.$
In this case, we define the stable holonomy $H_{\omega,\omega^{\prime}}^{s}$ by
$H_{\omega,\omega^{\prime}}^{s}=H_{\omega,\omega^{\prime}}^{s,n_{0}}.$
The unstable holonomy $H_{\omega,\omega^{\prime}}^{u}$ for $\omega^{\prime}\in W_{\rm loc}^{u}(\omega)$ is defined similarly by
$H_{\omega,\omega^{\prime}}^{u}=\bigl[A^{E}_{n}(\omega^{\prime})\bigr]^{-1}A^{E}_{n}(\omega)\qquad\text{for all}\quad n\leqslant-n_{0}.$
These abstract definitions of holonomies work not only for the cocycle (2.5), but also for any locally constant function $A^{E}:\Omega\to{\rm SL}(2,{\mathbb{R}})$. However, if $A^{E}$ is defined by (2.5), then the matrices $H_{\omega,\omega^{\prime}}^{s}$ and $H_{\omega,\omega^{\prime}}^{u}$ become very specific.
Proposition 2.4.
Let $A^{E}$ be defined in (2.5). Then
$H_{\omega,\omega^{\prime}}^{s}=\bigl[A^{E}(\omega^{\prime})\bigr]^{-1}A^{E}(\omega),\qquad\text{for any}\quad\omega^{\prime}\in W^{s}(\omega).$ (2.6)
Similarly,
$H_{\omega,\omega^{\prime}}^{u}={\rm Id},\qquad\text{for any}\quad\omega^{\prime}\in W^{u}(\omega).$ (2.7)
The general theory of dynamical systems tells us that the cocycle
$(T,A^{E}):\Omega\times{\mathbb{R}\mathbb{P}}^{1}\to{\mathbb{R}\mathbb{P}}^{1}$
defined by
$(T,A^{E})(\omega,\xi)=(T\omega,A^{E}(\omega)\xi)$
has an invariant probability measure $m$ on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}.$ We say that such a measure $m$ projects to $\mu$ if $m(\Delta\times{\mathbb{R}\mathbb{P}}^{1})=\mu(\Delta)$ for all Borel subsets $\Delta$ of $\Omega$. Given any $T$-invariant measure $\mu$ on $\Omega$, one can find a $(T,A^{E})$-invariant measure $m$ that projects to $\mu$ by applying the standard Krylov-Bogolyubov trick used to construct invariant measures.
Definition. Suppose $m$ is a $(T,A^{E})$-invariant probability measure on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}$ that projects to $\mu$. A disintegration of $m$ is a measurable family $\{m_{\omega}:\quad\omega\in\Omega\}$ of probability measures on ${\mathbb{R}\mathbb{P}}^{1}$ having the property
$m(D)=\int_{\Omega}m_{\omega}(\{\xi\in{\mathbb{R}\mathbb{P}}^{1}:\,\,(\omega,\xi)\in D\})d\mu(\omega)$
for each measurable set $D\subset\Omega\times{\mathbb{R}\mathbb{P}}^{1}.$
Existence of such a disintegration is guaranteed by Rokhlin’s theorem. Moreover, if $\{\tilde{m}_{\omega}:\quad\omega\in\Omega\}$ is another disintegration of $m$ then $m_{\omega}=\tilde{m}_{\omega}$ for $\mu$-almost every $\omega\in\Omega$. It is easy to see that $m$ is $(T,A^{E})$-invariant if and only if $A^{E}(\omega)_{*}m_{\omega}=m_{T\omega}$ for $\mu$-almost every $\omega\in\Omega$.
Definition. A $(T,A^{E})$-invariant measure $m$ on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}$ that projects to $\mu$ is said to be an s-state for $A^{E}$ provided it has a disintegration $\{m_{\omega}:\quad\omega\in\Omega\}$ such that for $\mu$-almost every $\omega\in\Omega$,
1)
$\quad A^{E}(\omega)_{*}m_{\omega}=m_{T\omega},$
2)
$\bigl(H^{s}_{\omega,\omega^{\prime}}\bigr)_{*}m_{\omega}=m_{\omega^{\prime}}\qquad\text{ for every}\quad\omega^{\prime}\in W^{s}(\omega).$
Definition. A $(T,A^{E})$-invariant measure $m$ on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}$ that projects to $\mu$ is said to be a u-state for $A^{E}$ provided it has a disintegration $\{m_{\omega}:\quad\omega\in\Omega\}$ such that for $\mu$-almost every $\omega\in\Omega$,
1)
$\quad A^{E}(\omega)_{*}m_{\omega}=m_{T\omega},$
2)
$\bigl(H^{u}_{\omega,\omega^{\prime}}\bigr)_{*}m_{\omega}=m_{\omega^{\prime}}\qquad\text{ for every}\quad\omega^{\prime}\in W^{u}(\omega)$
Let $\mathcal{E}$ be the set of energies $E\geqslant 0$ for which there is a measure $m$ that is both a u-state and an s-state. It was shown in [27] that the set $\mathcal{E}$ is discrete. It is convenient to add the squares of integer multiples of $\pi$ to $\mathcal{F}$, and consider the union
$\tilde{\mathcal{E}}={\mathcal{E}}\cup\bigcup_{n=0}^{\infty}\{(\pi n)^{2}\}.$
For $\eta>0$, we define the set
$B_{\eta}(\tilde{\mathcal{E}})=\{E>0:\,\,{\rm dist}(E,\tilde{\mathcal{E}})<\eta\}.$
[…]
Lemma 2.5.
For every $E\in I$ the cocycle $A^{E}$ has a unique u-state $m^{u,E}$. It depends on $E\in I$ continuously in the weak-* topology.
The proof of this lemma is absolutely the same as that of Lemma 3.4 in [2].
Below, we use the notation $\Omega^{\pm}_{j}=\pi^{\pm}[0;j]$. For a point $\omega^{-,j}\in\Omega_{j}^{-}$, we define $W^{u}_{\rm loc}(\omega^{-,j})=W^{u}_{\rm loc}(\omega)$ where $\omega$ is any point in $\Omega$ with the property $\pi^{-}\omega=\omega^{-,j}$.
The following lemma was proved in [2].
Lemma 2.6.
For a fixed $\omega^{-,j}\in\Omega_{j}^{-}$, let $\nu^{u}$ be a probability measure on $W^{u}_{\rm loc}(\omega^{-,j})$ having the property that
$C^{-1}\leqslant\frac{d(\pi_{*}^{+}\nu^{u})}{d\mu_{j}^{+}}\leqslant C.$
Then
$\frac{1}{n}\sum_{k=0}^{n-1}T^{k}_{*}\nu^{u}\to\mu,\quad\text{as}\quad n\to\infty,$
in the weak -* topology uniformly in $\omega^{-,j}$, $1\leqslant j\leqslant\ell$ and $\nu$.
Note that the notion of uniform convergence for a sequence such measures makes sense, because the collection of the measures that we consider is a metrizable space.
In what follows, we use the notation $F^{E}$ for the mapping from $\Omega\times\mathbb{RP}^{1}$ to itself defined by
$F^{E}(\omega,v)=(T\omega,A^{E}(\omega)v).$
Lemma 2.7.
Let $E\in I$ and $\omega^{-,j}\in\Omega_{j}^{-}$. Suppose $m$ is a probability measure on $W^{u}_{loc}(\omega^{-,j})\times{\mathbb{R}\mathbb{P}}^{1}$, whose projection $\nu^{u}$ to $W^{u}_{loc}(\omega^{-,j})$ satisfies the assumptions of Lemma 2.6. Then
$\frac{1}{n}\sum_{k=0}^{n-1}(F^{E})^{k}_{*}m\to m^{u,E}$ (2.8)
in the weak-* topology, uniformly in $\omega^{-,j}$, $E\in I$, and such choices of $m$.
Proof. By compactness of $\Omega\times{\mathbb{R}\mathbb{P}}^{1}$, we see that the collection of measures of the form
$\frac{1}{n}\sum_{k=0}^{n-1}(F^{E})^{k}_{*}m$ (2.9)
has weak-* accumulation points. Consider such an accumulation point and denote it by $\tilde{m}$. Without loss of generality, we may just assume that
$\frac{1}{n}\sum_{k=0}^{n-1}(F^{E})^{k}_{*}m\to\tilde{m}$. (2.10)
We only need to prove that $\tilde{m}=m^{u,E}$ and that the convergence is uniform. Clearly, $\tilde{m}$ is invariant under $F^{E}$ and it projects to the limit of the sequence
$\frac{1}{n}\sum_{k=0}^{n-1}T^{k}_{*}\nu^{u}$ (2.11)
in the first component. By Lemma 2.6, it means that $\tilde{m}$ projects to $\mu$.
Let us show now that any disintegration $\{\tilde{m}_{\omega}\}_{\omega\in\Omega}$ of $\tilde{m}$ is invariant under the unstable holonomy. Since $A$ depends only on the past, any unstable holonomy is the identity $\rm Id$}. Thus we only need to show that $\tilde{m}_{\omega}=\tilde{m}_{\omega^{\prime}}$ for $\omega$ and $\omega^{\prime}$ in the same local unstable set.
For this purpose, we let $\tilde{m}^{-}=(\pi^{-}\times{\rm Id})_{*}\tilde{m}$, which is a measure on $\Omega_{-}\times{\mathbb{R}\mathbb{P}}^{1}$.
Since $A^{E}$ depends only on the past, it naturally descends to a map on $\Omega_{-}$. Namely, $A^{E}(\omega)=\tilde{A}^{E}(\pi^{-}\omega)$ for some map $\tilde{A}^{E}:\Omega_{-}\to{\rm SL}(2,{\mathbb{R}})$, which we will still denote by $A^{E}$.
Let $F^{E}_{-}$ be the action of $(T_{-},A_{-1}^{E})$ on $\Omega_{-}\times{\mathbb{R}P}^{1}$ where $T_{-}$ is the right shift on $\Omega_{-}$ and $A_{-1}^{E}(\omega_{-})=A^{E}(T_{-}\omega_{-})$. Then
$(\pi^{-}\times{\rm Id})\circ(T,A^{E})^{-1}=(T_{-},A_{-1}^{E})\circ(\pi^{-}\times{\rm Id}),$ (2.12)
which may be written in the form $(\pi_{-}\times{\rm Id})\circ(F^{E})^{-1}=F^{E}_{-}(\pi_{-}\times{\rm Id}).$ This implies that $\tilde{m}^{-}$ is invariant under $F^{E}_{-}$. Indeed,
[…]
Let $\{\tilde{m}^{-}_{\omega_{-}}\}$ be a disintegration of $\tilde{m}^{-}$. By $F^{E}_{-}$-invariance, we then have $A^{E}(T_{-}\omega_{-})]^{-1}_{*}\tilde{m}^{-}_{\omega_{-}}=\tilde{m}^{-}_{T_{-}\omega_{-}}$, or equivalently,
$(A^{E}_{-1}(\omega_{-}))_{*}\tilde{m}^{-}_{\omega_{-}}=\tilde{m}^{-}_{T_{-}\omega_{-}}.$ (2.14)
By a special case of [3], Lemma 3.4, the disintegration $\{\tilde{m}_{\omega}\}$ of $\tilde{m}$ can be recovered from the disintegration $\{\tilde{m}^{-}_{\omega_{-}}\}$ of $\tilde{m}^{-}$ via
$\tilde{m}_{\omega}=\lim_{n\to\infty}(A^{E}_{-n}(\pi_{-}(T^{n}\omega)))_{*}\tilde{m}^{-}_{\pi^{-}(T^{n}\omega)}$ (2.15)
This implies that
$\tilde{m}_{\omega}=\tilde{m}^{-}_{\pi^{-}\omega}$ (2.16)
so $\tilde{m}_{\omega}$ is constant on the local unstable set. This concludes the proof that $\tilde{m}$ is a u-state. Therefore, by uniqueness of the u-state, it must be equal to $m^{u,E}$. Uniform convergence follows again from uniqueness of the limit. Indeed, let $\rho$ be the metric on the space of probability measures defined on $\Omega\times\mathbb{RP}^{1}$. For instance, such a metric may be chosen as
$\rho(m,m^{\prime})=\sum_{s=1}^{\infty}2^{-s}\Bigl|\int_{\Omega\times\mathbb{RP}^{1}}f_{s}dm-\int_{\Omega\times\mathbb{RP}^{1}}f_{s}dm^{\prime}\Bigr|,$ (2.17)
where $\{f_{s}\}_{s=1}^{\infty}$ is a dense subset of the unit ball in $C(\Omega\times\mathbb{RP}^{1})$.
Let $\varepsilon>0$. Suppose there is a sequence of measures $m_{j}$ satisfying conditions of the lemma and a sequence of energies $E_{j}\to E$ such that
$\rho\Bigl(\frac{1}{n_{j}}\sum_{k=0}^{n_{j}-1}(F^{E_{j}})^{k}_{*}m_{j},\,m^{u,E_{j},}\Bigr)>\varepsilon\qquad\forall j.$ (2.18)
By compactness, we may assume that $\frac{1}{n_{j}}\sum_{k=0}^{n_{j}-1}(F^{E_{j}})^{k}_{*}m_{j}$ converges to some probability measure $\tilde{m}$. But then
$\frac{1}{n_{j}}\sum_{k=0}^{n_{j}-1}(F^{E_{j}})^{k+1}_{*}m_{j}-\frac{1}{n_{j}}\sum_{k=0}^{n_{j}-1}(F^{E_{j}})^{k}_{*}m_{j}\to 0,\qquad\text{as}\quad j\to\infty$ (2.19)
implies that the sequence $F^{E_{j}}\tilde{m}$ also converges to $\tilde{m}$. Thus, the measure $\tilde{m}$ is $F^{E}$-invariant. Repeating the arguments that lead us to (2.16), we conclude that $\tilde{m}$ is a u-state for $A^{E}$. On the other hand, (2.18) implies that
$\rho\Bigl(\tilde{m},\,m^{u,E}\Bigr)\geqslant\varepsilon,$
which contradicts the fact that $\tilde{m}=m^{u,E}$. $\Box$
A disintegration of $\mu$ with respect to the local unstable sets is a (measurable) family of measures $\mu^{u}_{\omega^{-}}$
$\{\mu^{u}_{\omega^{-}}:\,\,\text{probability measure on}\,W^{u}_{\rm loc}(\omega^{-})\}_{\omega^{-}\in\Omega_{-}}$
having the property that
$\mu(D)=\int_{\Omega_{-}}\mu^{u}_{\omega^{-}}\{\omega\in D:\,\,\pi^{-}(\omega)=\omega^{-}\}\,d\mu^{-}(\omega^{-})$
Such a disintegration exists by Rokhlin’s theorem.
For a continuous function $\phi$ on $\Omega\times{\mathbb{R}\mathbb{P}}^{1}$, define $S_{n}(\phi)$ by
$S_{n}(\phi)=\frac{1}{n}\sum_{k=0}^{n-1}\phi\circ(F^{E})^{k}.$
Theorem 2.8.
For every $\phi=\bar{\phi}\in C^{\alpha}(\Omega\times{\mathbb{R}\mathbb{P}}^{1})$, $\varepsilon>0$, and $E_{0}\in I$, there are positive constants $C$, $c$, and $r$ such that
$\mu^{u}_{\omega^{-}}\Bigl\{\omega\in W^{u}_{\rm loc}(\omega^{-}):\,\,\Bigl|\frac{1}{n}S_{n}(\phi)(\omega,v)-\int\phi\,dm^{u,E}\Bigr|>\varepsilon\Bigr\}<Ce^{cn}$
uniformly in $(\omega^{-},v)\in\Omega_{-}\times{\mathbb{R}\mathbb{P}}^{1}$ and $E\in I\cap(E_{0}-r,E_{0}+r)$.
The proof of this theorem is based on the following lemma stated in [2].
Lemma 2.9.
Let $\{d_{n}\}_{n\in{\mathbb{N}}}$ be a sequence of random variables such that
${\mathbb{E}}(d_{n+1}|{\mathfrak{F}}_{n})=0,$
where ${\mathfrak{F}}_{n}$ is the $\sigma$-algebra generated by the functions $d_{1},\dots,d_{n}$. Suppose that
$\|d_{n}\|_{\infty}\leqslant a,\qquad\forall n.$
Then for every $\varepsilon>0$, there is a constant $c>0$ such that
${\mathbb{P}}\Bigl(\Bigl|\frac{1}{N}\sum_{n=1}^{N}d_{n}\Bigr|>\varepsilon\Bigr)<e^{-\frac{c\varepsilon^{2}N}{2a^{2}}},\qquad\forall N\geqslant 1.$
Proof of Theorem  2.8. While we have made minor cosmetic changes to the proof provided by the authors of [2], the core ideas remain the same. Given its elegance, we felt it necessary to include this proof.
By the bounded distortion property of $\mu$, there exists a constant $C\geqslant 1$ such that for each $1\leqslant j\leqslant\ell$ and $\mu^{-}$-almost every $\omega^{-,j}\in\Omega_{j}^{-}$
$C^{-1}\leqslant\frac{d(\pi^{+}_{*}\mu^{u}_{\omega^{-,j}})}{d\mu^{+}_{j}}\leqslant C.$ (2.20)
Indeed, let $\omega\in\Omega_{j}^{+}$, then
$\begin{split}\frac{d(\pi^{+}_{*}\mu^{u}_{\omega^{-,j}})}{d\mu^{+}_{j}}(\omega)=\lim_{n\to\infty}\frac{\pi^{+}_{*}\mu^{u}_{\omega^{-,j}}([0,\omega_{0},\omega_{1},\dots,\omega_{n}])}{\mu^{+}_{j}([0,\omega_{0},\omega_{1},\dots,\omega_{n}])}=\\ \lim_{n\to\infty}\frac{\mu^{u}_{\omega^{-,j}}(W^{u}_{\rm loc}(\omega^{-,j})\cap[0,\omega_{0},\omega_{1},\dots,\omega_{n}])}{\mu^{+}_{j}([0,\omega_{0},\omega_{1},\dots,\omega_{n}])}=\\ \lim_{n\to\infty}\lim_{l\to\infty}\frac{\int_{[-l;i_{-l},\dots,i_{-1},j]}\mu^{u}_{\omega^{-,j}}(W^{u}_{\rm loc}(\omega^{-,j})\cap[0,\omega_{0},\omega_{1},\dots,\omega_{n}])d\omega^{-,j}}{\mu_{j}^{-}([-l;i_{-l},\dots,i_{-1}])\mu^{+}_{j}([0,\omega_{0},\omega_{1},\dots,\omega_{n}])}=\\ \lim_{n\to\infty}\lim_{l\to\infty}\frac{\mu([-l;i_{-l},\dots,i_{-1},]\cap[0,\omega_{0},\omega_{1},\dots,\omega_{n}])}{\mu_{j}^{-}([-l;i_{-l},\dots,i_{-1}])\mu^{+}_{j}([0,\omega_{0},\omega_{1},\dots,\omega_{n}])}.\end{split}$
Without loss of generality, we will assume that inequality (2.20) holds not only almost everywhere, but for all $\omega^{-,j}\in\Omega^{-,j}$, and, therefore, $\mu^{u}_{\omega^{-,j}}$ satisfies the assumption of Lemma 2.6.
We will work with integrals of a lift of the measure $\mu^{u}_{\omega^{j,-}}$ defined as follows. Fix any $v\in{\mathbb{R}\mathbb{P}}^{1}$ and set $m$ to be the product of the measures $\mu^{u}_{\omega^{j,-}}$ and $\delta_{v}$ defined on $W_{\rm loc}^{u}(\omega^{j,-})\times{\mathbb{R}\mathbb{P}}^{1}$. Then $m$ satisfies the assumptions of Lemma  2.7.
Since $\pi^{+}T=T_{+}\pi^{+}$, we have the relations $\pi^{+}T^{s}=(T_{+})^{s}\pi^{+}$, and $\pi^{+}_{*}T^{s}_{*}\mu^{u}_{\omega^{-,j}}=(T_{+})^{s}_{*}\pi^{+}_{*}\mu^{u}_{\omega^{-,j}}$, which imply that
$C^{-1}\leqslant\frac{d(\pi^{+}_{*}T^{s}_{*}\mu^{u}_{\omega^{-,j}})}{d(T_{+})^{s}_{*}\mu^{+}_{j}}\leqslant C.$
By $T_{+}$-invariance of $\mu^{+}$, and the bounded distortion property of the measure, for admissible $j,i_{1},\dots,i_{s}$, we have
$C^{-1}\mu([0;j,i_{1},\dots,i_{s-1}])\leqslant\frac{d((T_{+})^{s}_{*}\mu^{+}_{j}|_{[0;j,i_{1},\dots,i_{s}]})}{d\mu_{i_{s}}^{+}}\leqslant C\mu([0;j,i_{1},\dots,i_{s-1}]).$
The above estimate implies that
$\frac{1}{\mu^{u}_{\omega^{-,j}([0;j,i_{1},\dots,i_{s}])}}T^{s}_{*}\bigl(\mu^{u}_{\omega^{-,j}}|_{[1;i_{1},\dots,,i_{s}]}\bigr)$
is a probability measure that satisfies (2.20), and thereby, the assumptions of Lemma 2.6. Moreover, the measure
$\frac{1}{m([0:j,i_{1},...,i_{s}]]\times{\mathbb{R}P}^{1})}(F^{E})^{s}_{*}\bigl(m|_{[1:i_{1},...,i_{s}]\times{\mathbb{R}\mathbb{P}}^{1}}\bigr)$ (2.21)
obeys the conditions of Lemma  2.7. Now for each $\omega\in W^{u}_{\rm loc}(\omega^{-,j})$, and $i\in{\mathbb{N}}$ we define the set
$D_{i}(\omega):=[1;\omega_{1},...,\omega_{i}]\cap W^{u}_{\rm loc}(\omega^{-,j})\times{\mathbb{R}\mathbb{P}}^{1}.$
If $\omega_{1}=i_{1},\dots,\omega_{s}=i_{s}$, then the measure (2.21) can be written as
$\frac{1}{m(D_{s}(\omega))}(F^{E})^{s}_{*}\bigl(m|_{D_{s}(\omega)}\bigr)$
Note also that for every continuous function $\psi$ on $\Omega$, we have
[…]
$\Bigl|\frac{1}{m(D_{i}(\omega))}\int_{D_{i}(\omega)}\frac{1}{N}S^{E}_{N}(\phi\circ(F^{E})^{i})dm-\int_{\Omega}\phi\,dm^{u,E}\Bigr|<\varepsilon/4$ (2.22)
Define now the functions $Y_{i}:W^{u}_{\rm loc}(\omega^{-,j})\to{\mathbb{R}}$ by setting
$Y_{i}(\omega)=\frac{1}{m(D_{i}(\omega))}\int_{D_{i}(\omega)}S^{E}_{i}(\phi)dm.$ (2.23)
Obviously, $Y_{i}$ depends only on $\omega_{0},\dots,\omega_{i}$. Let ${\mathfrak{B}}_{i}$ be the $\sigma$-algebra generated by the functions $Y_{0},\dots,Y_{i}$, which is basically generated by the cylinder sets $0;n_{0},\dots,n_{i}]$. In particular, the conditional expectation of $Y_{i+N}$ with respect to ${\mathfrak{B}}_{i}$ is
$\begin{split}{\mathbb{E}}\Bigl(Y_{i+N}|{\mathfrak{B}}_{i}\Bigr)(\omega)=\frac{1}{m(D_{i}(\omega))}\sum_{D_{i+N}(\tilde{\omega})\subset D_{i}(\omega)}m(D_{i+N}(\tilde{\omega}))Y_{i+N}(\tilde{\omega})\\ =\frac{1}{m(D_{i}(\omega))}\sum_{D_{i+N}(\tilde{\omega})\subset D_{i}(\omega)}\int_{D_{i+N}(\tilde{\omega})}S^{E}_{i+N}(\phi)\,dm\\ =\frac{1}{m(D_{i}(\omega))}\int_{D_{i}(\omega)}S^{E}_{i+N}(\phi)\,dm\end{split}$ (2.24)
Clearly,
${\mathbb{E}}(Y_{i}|{\mathfrak{B}}_{i})=Y_{i}.$ (2.25)
Thus, the estimate (2.22) can be rewritten as follows
$\Bigl|\frac{1}{N}{\mathbb{E}}\Bigl(Y_{N+i}-Y_{i}|{\mathfrak{B}}_{i}\Bigr)-\int_{\Omega}\phi dm^{u,E}\Bigr|<\varepsilon/4.$ (2.26)
Define now the sequence
$X_{n}=Y_{nN}-\sum_{k=1}^{n}{\mathbb{E}}\Bigl(Y_{kN}-Y_{(k-1)N}|{\mathfrak{B}}_{(k-1)N}\Bigr).$ (2.27)
It is easy to see that the sequence $\{X_{n}\}$ is a martingale, that is, (LABEL:21) holds for $d_{n}=X_{n+1}-X_{n}$. Indeed,
$\begin{split}X_{n+1}-X_{n}=Y_{(n+1)N}-Y_{nN}-{\mathbb{E}}\Bigl(Y_{(n+1)N}-Y_{nN}|{\mathfrak{B}}_{nN}\Bigr)=\\ Y_{(n+1)N}-{\mathbb{E}}\Bigl(Y_{(n+1)N}|{\mathfrak{B}}_{nN}\Bigr).\end{split}$ (2.28)
Since the $\sigma$-algebra ${\mathfrak{F}}_{n}$ generated by the functions $X_{1},\dots,X_{2}$ is precisely ${\mathfrak{B}}_{nN}$, the above equation implies that for all $n\geqslant 1$,
${\mathbb{E}}\Bigl(X_{n+1}-X_{n}|_{{\mathfrak{F}}_{n}}\Bigr)=0$ (2.29)
We claim that
$\sup_{n}\|X_{n+1}-X_{n}\|_{\infty}\leqslant a<\infty,$ (2.30)
because of the Hölder continuity of $\phi$. Indeed, it is clear that
$|X_{1}|=|Y_{N}-{\mathbb{E}}(Y_{N}-Y_{0}|{\mathfrak{B}}_{0})|\leqslant(2\|\phi\|_{\infty}+1)N=CN.$
On the other hand, $X_{n+1}-X_{n}$ may be rewritten as
$\frac{1}{m(D_{(n+1)N}(\omega))}\int_{D_{(n+1)N}(\omega)}S_{(n+1)N}^{E}(\phi)dm-\frac{1}{m(D_{Nn}(\omega))}\int_{D_{Nn}(\omega)}S_{(n+1)N}^{E}(\phi)dm$ (2.31)
Clearly, it is sufficient to prove that the difference of two values of $S_{(N+1)n}^{E}(\phi)$ at the points $(\omega^{\prime},v)$ and $(\omega^{\prime\prime},v)$ does not exceed a constant, if $\omega^{\prime},\omega^{\prime\prime}\in D_{Nn}(\omega)$. Note that $A^{E}_{i}(\omega^{\prime})v$ is independent of $\omega$ for all $\omega^{\prime}\in D_{nN}(\omega)$ and all $0\leqslant i\leqslant nN$. So, we may denote $A^{E}_{i}(\omega^{\prime})v$ by $v_{i}$. Now, for any $\omega^{\prime},\omega^{\prime\prime}\in D_{nN}(\omega)$,
$\begin{split}|S_{(n+1)N}^{E}(\phi)(\omega^{\prime},v)-S_{(n+1)N}^{E}(\phi)(\omega^{\prime\prime},v)|\leqslant\\ \Bigl(\sum_{i=0}^{nN-1}+\sum_{i=nN}^{(n+1)N-1}\Bigr)|\phi((F^{E})^{i}(\omega^{\prime},v))-\phi((F^{E})^{i}(\omega^{\prime},v))|\leqslant\\ \sum_{i=0}^{nN-1}|\phi(T^{i}\omega^{\prime},v_{i})-\phi(T^{i}\omega^{\prime\prime},v_{i})|+CN\leqslant C\sum_{i=0}^{nN-1}d(T^{i}\omega^{\prime},T^{i}\omega^{\prime\prime})^{\alpha}+CN\\ \leqslant C\sum_{i=0}^{nN-1}e^{-\alpha(nN-i)}+CN\leqslant C_{\phi}+CN.\end{split}$ (2.32)
Thus, it follows from Lemma  2.9 that for every $\delta>0$ and all $n\geqslant 1$, we have
$\mu^{u}_{\omega^{-,j}}\Bigl\{\omega\in W^{u}_{\rm loc}(\omega^{-,j}):\,\frac{1}{n}X_{n}>\delta\Bigr\}={\mathbb{P}}\Bigl\{\frac{1}{n}X_{n}>\delta\Bigr\}<e^{-\frac{\delta^{2}}{2a^{2}}n}$ (2.33)
Suppose $\omega\in W^{u}_{\rm loc}(\omega^{-,j})$ satisfies
$\Bigl|\frac{1}{nN}S_{nN}^{E}(\phi)(n,v)-\int_{\Omega\times{\mathbb{R}\mathbb{P}}^{1}}\phi\,dm^{u,E}\Bigr|>\varepsilon$ (2.34)
We claim that this inequality holds on the set $D_{nN}(\omega)$ with $\varepsilon/2$. Indeed, repeating the arguments that worked for the bound of $\|X_{n+1}-X_{n}\|_{\infty}$, we conclude that for all $\omega^{\prime}\in D_{nN}(\omega)$ $0,\omega_{0},\omega_{1},\dots,\omega_{nN}]\cap W^{u}_{\rm loc}(\omega^{-,j})$,
$|S_{nN}^{E}(\phi)(\omega^{\prime},v)-S_{nN}^{E}(\phi)(\omega,v)|<C_{\phi}.$
Combining this estimate with (2.34), we obtain that
$\Bigl|\frac{1}{nN}Y_{nN}(\omega)-\int_{\Omega\times{\mathbb{R}\mathbb{P}}^{1}}\phi\,dm^{u,E}\Bigr|>\varepsilon/2$ (2.35)
for all $n\geqslant N_{0}$, where $N_{0}$ depends only on $\phi$ and $\varepsilon$. Combining the latter inequality with (2.26), we conclude that if $\omega$ satisfies (2.34), then
$\begin{split}\Bigl|\frac{1}{nN}X_{nN}\Bigr|=\Bigl|\frac{1}{nN}Y_{nN}-\frac{1}{n}\sum_{k=1}^{n}\frac{1}{N}{\mathbb{E}}\Bigl(Y_{kN}-Y_{(k-1)N}|_{{\mathfrak{B}}_{(k-1)N}}\Bigr)\Bigr|\\ \geqslant\Bigl|\frac{1}{nN}Y_{nN}-\int_{\Omega\times{\mathbb{R}\mathbb{P}}^{1}}\phi\,dm^{u,E}\Bigr|-\\ \Bigl|\frac{1}{n}\sum_{k=1}^{n}\Bigl(\frac{1}{N}{\mathbb{E}}\bigl(Y_{kN}-Y_{(k-1)N}|_{{\mathfrak{B}}_{(k-1)N}}\bigr)-\int_{\Omega\times{\mathbb{R}\mathbb{P}}^{1}}\phi\,dm^{u,E}\Bigr)\Bigr|>\varepsilon/4.\end{split}$ (2.36)
Applying (2.33) with $\delta=\varepsilon N/4$, we obtain for all $n\geqslant N_{0}$,
$\mu^{u}_{\omega^{-,j}}\Bigl\{\omega\in W^{u}_{\rm loc}(\omega^{-,j}):\,\,\Bigl|\frac{1}{nN}S_{nN}^{E}(\phi)(\omega,v)-\int_{\Omega\times{\mathbb{R}\mathbb{P}}^{1}}\phi\,dm^{u,E}\Bigr|>\varepsilon\Bigr\}<e^{-\frac{\varepsilon^{2}N}{32a^{2}}nN}.$ (2.37)
It remains to note that the same arguments could be repeated for a sequence of indices $nN+l$ replacing the sequence $nN$. The only property of this sequence that we used is that the distance between two consecutive indices in the sequence is $N$. This completes the proof of Theorem 2.8. $\Box$
For $v\in{\mathbb{R}\mathbb{P}}^{1}$, we define $\underline{v}$ as the unit vector having the same direction as $v$. It is shown in [2] that Theorem 2.8 implies the following result.
Theorem 2.10.
For any $\varepsilon>0$, there are positive constants $C$ and $c$ for which
$\mu^{u}_{\omega^{-}}\Bigl\{\omega\in W^{u}_{\rm loc}(\omega^{-}):\,\,\bigl|\frac{1}{n}\ln\|A^{E}_{n}(\omega)\underline{v}\|-L(E)\bigr|>\varepsilon\Bigr\}<Ce^{-cn}$
uniformly in $(\omega^{-},v)\in\Omega_{-}\times{\mathbb{R}\mathbb{P}}^{1}$, and $E\in I$.
It is relatively easy to establish (see [2] again) that Theorem 2.10 implies Theorem 2.11 below.
Theorem 2.11.
For any $\varepsilon>0$, there are positive constants $C$ and $c$ depending only on $\varepsilon$ for which
$\mu\Bigl\{\omega\in\Omega:\,\,\bigl|\frac{1}{n}\ln\|A^{E}_{n}(\omega)\|-L(E)\bigr|>\varepsilon\Bigr\}<Ce^{-cn}$
uniformly in $E\in I$.
We will call this theorem the Uniform Large Deviations (ULD).
3. Localization as a consequence of ULD
It is convenient for our purposes to use the fillowing lemma from [2].
Lemma 3.1.
Let $F$ be an $\alpha$-Hölder continuous function on $\Omega_{+}$ obeying the condition $\|F\|_{\infty}<1$ and $|F(\omega)-F(\omega^{\prime})|<Kd(\omega,\omega^{\prime})^{\alpha}$ with $K>1$. Then for all $\varepsilon>0$ and all $r,n\geqslant 1$ we have
$\mu^{+}\Bigl\{\omega^{+}:\,\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}F(T_{+}^{ns}\omega)-\int Fd\mu^{+}\Big|>\varepsilon\Bigr\}<\exp\Bigl(-\frac{c\varepsilon^{2}n^{2}r}{\log^{2}(K\varepsilon^{-1})}\Bigr)$ (3.1)
where the constant $c>0$ is independent of $n,r,K$ and $\varepsilon$.
In what follows, we will use the notations
$g_{n}(\omega,E)=\frac{1}{n}\log(\|A_{n}^{E}(\omega)\|),$
and $\Gamma=\sup_{E\in I}\|A^{E}(\cdot)\|_{\infty}$. For each $1\leqslant j\leqslant\ell$, choose $\omega^{(j)}\in[0,j]$ and define $\phi(\omega)=\omega^{(\omega_{0})}\wedge\omega$ to be the unique element that belongs to $W^{u}_{\rm loc}(\omega^{(\omega_{0})})\cap W^{s}_{\rm loc}(\omega)$.
For each natural number $n$, we define the function
$g^{+}_{n}(\omega,E)=g_{n-1}(T\omega,E),\qquad\omega\in\Omega.$ (3.2)
Then we set
$\begin{split}h_{n}(\omega,E)=g_{n}(\omega,E)-g^{+}_{n}(\omega,E),\qquad\omega\in\Omega.\end{split}$ (3.3)
Note that
$\sup_{E\in I}\|g_{n}^{+}(\cdot,E)\|_{\infty}=\sup_{E\in I}\|g_{n}(\cdot,E)\|_{\infty}\leqslant\ln\Gamma.$ (3.4)
Obviously $g^{+}_{n}(\omega,E)$, it is constant on $W^{s}_{\rm loc}(\omega)$ (it depends only on the future). Moreover, it depends only on $\omega_{0},\dots,\omega_{n-1}$. Therefore,
$|g^{+}_{n}(\omega,E)-g^{+}_{n}(\omega^{\prime},E)|\leqslant 2e^{n\alpha}\ln\Gamma\cdot d(\omega,\omega^{\prime})^{\alpha},$ (3.5)
for any $\alpha>0$.
Proposition 3.2.
Let $h_{n}$ be defined by (3.3), then
$|h_{n}(\omega,E)|\leqslant\frac{2}{n}\bigl(\ln(\Gamma)+1\bigr).$ (3.6)
Proof. Note that
$\ln\|A_{n-1}^{E}(T\omega)\|-\ln\|[A^{E}(\omega)]^{-1}\|\leqslant\ln\|A_{n}^{E}(\omega)\|\leqslant\ln\|A_{n-1}^{E}(T\omega)\|+\ln\|A^{E}(\omega)\|.$
Consequently,
$\Bigl|g_{n}(\omega,E)-\frac{n-1}{n}g_{n-1}(T\omega,E)\Bigr|\leqslant\frac{1}{n}\bigl(\ln(\Gamma)+1\bigr).$
Finally, since $g_{n-1}(T\omega,E)$ and $g^{+}_{n}(\omega),E)$ are the same, we obtain (3.6). $\Box$
In the statement below, we treat $g_{n}^{+}(\cdot,E)$ as a function defined on $\Omega^{+}$.
Lemma $3.$
For every $\varepsilon\in(0,1)$, there is an $n_{0}=n_{0}(\varepsilon,\Gamma)$ such that
$\mu^{+}\Bigl\{\omega^{+}:\,\,\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g^{+}_{n}(T_{+}^{ns}\omega^{+},E)-L(E)\Bigr|>\varepsilon\Bigr\}\leqslant e^{-c\varepsilon^{2}r}$
for all $E\in I,r\in{\mathbb{N}}$, and all $n\geqslant n_{0}$. The constant $c>0$ in this inequality is universal.
Proof. It is easy to see that the ULD property of $g_{n}$ imples a similar property for $g_{n}^{+}$. That is, for every $\varepsilon>0$, there are positive constants $C$ and $c$ such that
$\mu^{+}\Bigl\{\omega^{+}:\,\,\Bigl|g_{n}(\omega^{+},E)-L(E)\Bigr|>\varepsilon\Bigr\}\leqslant Ce^{-cn}$
holds uniformly in $E\in I$. Since the function $g_{n}^{+}$ is bounded by $\ln\Gamma$ and $0<L(E)\leqslant\ln\Gamma$, we conclude that there is an $n_{1}(\varepsilon)$ such that
$\Bigl|\int_{\Omega^{+}}g_{n}(\omega^{+},E)d\mu^{+}-L(E)\Bigr|\leqslant\frac{\varepsilon}{10}$
for all $n>n_{1}(\varepsilon)$. We apply Lemma 3.1 to the function $F=\frac{g_{n}^{+}}{2\ln\Gamma}.$ Setting $K=3e^{n\alpha}$, we obtain that there is an $n_{2}(\varepsilon,\Gamma)$ such that
$\mu^{+}\Bigl\{\omega^{+}:\,\,\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g^{+}_{n}(T_{+}^{ns}\omega^{+},E)-\int_{\Omega^{+}}g^{+}_{n}(\omega,E)d\mu^{+}\Bigr|>\varepsilon\Bigr\}\leqslant e^{-c\varepsilon^{2}r}$
for all $n>n_{2}(\varepsilon,\Gamma)$. It remains to combine the two last estimates. $\Box$
Lemma 3.4.
For every $\varepsilon\in(0,1)$, there is an $n_{0}=n_{0}(\varepsilon,\Gamma)$ such that
$\mu\Bigl\{\omega\in\Omega:\,\,\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g_{n}(T^{ns+s_{0}}\omega,E)-L(E)\Bigr|>\varepsilon\Bigr\}\leqslant e^{-c\varepsilon^{2}r},$ (3.7)
for all $n>n_{0}$, $s_{0}\in{\mathbb{Z}}$ and $r\in{\mathbb{N}}$. The constant $c>0$ in this inequality is universal.
Proof. Let
$B_{r}^{+}(\varepsilon)=\Bigl\{\omega^{+}\in\Omega^{+}:\,\,\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g^{+}_{n}(T_{+}^{ns}\omega^{+},E)-L(E)\Bigr|>\varepsilon\Bigr\}.$
Using Proposition 3.2, we observe that
$|g^{+}_{n}(\omega^{+},E)-g_{n}(\omega,E)|\leqslant\frac{2}{n}\bigl(\ln(\Gamma)+1\bigr).$ (3.8)
If $n>4\frac{(1+\ln\Gamma)}{\varepsilon}$, then
$\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g_{n}(T^{ns}\omega,E)-L(E)\Bigr|>\varepsilon\implies\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g^{+}_{n}(T_{+}^{ns}\omega^{+},E)-L(E)\Bigr|>\varepsilon/2.$
Consequently,
$\begin{split}\mu\Bigl\{\omega\in\Omega:\,\,\Bigl|\frac{1}{r}\sum_{s=0}^{r-1}g_{n}(T^{ns}\omega,E)-L(E)\Bigr|>\varepsilon\Bigr\}\leqslant\mu\Bigl((\pi^{+})^{-1}B_{r}^{+}(\varepsilon/2)\Bigr)\\ =\mu^{+}\Bigl(B_{r}^{+}(\varepsilon/2)\Bigr)\leqslant e^{-c\varepsilon^{2}r}.\end{split}$
Finally, one can add $s_{0}$ to $ns$, because $\mu$ is $T$-invariant. $\qquad\Box$
It is established in the last section of this paper that there are constants $\beta>0$ and $C>0$ such that
$|L(E^{\prime})-L(E)|\leqslant C|E^{\prime}-E|^{\beta},\qquad\forall E,E^{\prime}\in I.$ (3.9)
In the proposition below, instead of the exponent $2/3$ obtained in [2], we obtain a smaller exponent $2/7$, which is justified by our analysis of a special case.
Proposition 3.5.
For any $0<\varepsilon<1$, there is a set $\Omega(\varepsilon)\subset\Omega$ of full $\mu$-measure such that for any $\omega\in\Omega(\varepsilon)$ there is an index $\tilde{n}_{0}=\tilde{n}_{0}(\omega,\varepsilon)$ for which
$\Bigl|\frac{1}{n^{4}}\sum_{s=0}^{n^{4}-1}g_{n}(T^{ns+s_{0}}\omega,E)-L(E)\Bigr|<\varepsilon,$ (3.10)
if $n>\max\{\tilde{n}_{0},(\ln(|s_{0}|+1))^{2/7}\}$ and $E\in I$.
Proof. Let us consider the complement of the set of $\omega$’s for which (3.10) holds for all $E\in I$:
$B_{n,s_{0}}=\Bigl\{\omega\in\Omega:\,\,\sup_{E\in I}\Bigl|\frac{1}{n^{4}}\sum_{s=0}^{n^{4}-1}g_{n}(T^{ns+s_{0}}\omega,E)-L(E)\Bigr|\geqslant\varepsilon\Bigr\}.$ (3.11)
Let $\varkappa=|I|$. For any $0<\delta\leqslant\varkappa/2$, define the discrete set $I_{0}=I\cap(\delta{\mathbb{Z}})$. Then the cardinality of the set $I_{0}$ obeys ${\rm card}(I_{0})\leqslant\varkappa/\delta+1\leqslant 2\varkappa/\delta.$ Choose now $\delta=\varepsilon/(3\Gamma^{n})$. If necessary, choose $n$ so large that
$C\delta^{\beta}=C\bigl(\frac{\varepsilon}{3\Gamma^{n}}\bigr)^{\beta}<\frac{\varepsilon}{3}$ (3.12)
wher $C$ and $\beta$ are the same as in (3.9).
Combining (3.9) with the inequality
$|g_{n}(\omega,E)-g_{n}(\omega,E^{\prime})|\leqslant\Gamma^{n-1}|E-E^{\prime}|,$
we obtain
$B_{n,s_{0}}\subset\bigcup_{E\in I_{0}}\Bigl\{\omega\in\Omega:\,\,\Bigl|\frac{1}{n^{4}}\sum_{s=0}^{n^{4}-1}g_{n}(T^{ns+s_{0}}\omega,E)-L(E)\Bigr|\geqslant\frac{\varepsilon}{3}\Bigr\}.$
Taking $n>n_{0}(\varepsilon/3,\Gamma)$ and using the fact that ${\rm card}(I_{0})\leqslant\frac{6\varkappa\Gamma^{n}}{\varepsilon}$, we obtain by Lemma 3.4 that
$\mu\bigl(B_{n,s_{0}}\bigr)\leqslant\frac{6\varkappa\Gamma^{n}}{\varepsilon}e^{-c\varepsilon^{2}n^{4}}$ (3.13)
Using this estimate, we conclude that the measures of the sets
$B_{n}=\bigcup_{|s_{0}|\leqslant e^{n^{7/2}}}B_{n,s_{0}}$
obey the condition
$\sum_{n=n_{0}+1}^{\infty}\mu(B_{n})<\infty.$
By the Borel-Cantelly lemma, almost every point $\omega$ belongs to at most finitely many of the sets $B_{n}$. In other words, for almost every point $\omega$, there is a number $\tilde{n}_{0}(\omega,\varepsilon)$ such that $\omega\notin B_{n,s_{0}}$ for all $|s_{0}|\leqslant e^{n^{7/2}}$ if $n>\tilde{n}_{0}$. It remains to use the definition (3.11) of the set $B_{n,s_{0}}$. $\Box$
In what follows below, $E\geqslant 0$ and $\tilde{E}\in[-2,2]$ are related by $\tilde{E}=2\cos(\sqrt{E})$.
The obtained results will be used to estimate the finite volume Green’s function. Let $P_{\Lambda}$ be the orthogonal projection from $\ell^{2}({\mathbb{Z}})$ onto the set $\ell^{2}(\Lambda)$ where $\Lambda=\{0,1,\dots,N-1\}$. Define the finite rank operator
${\mathcal{H}}_{\omega,N}=P_{\Lambda}{\mathcal{H}}_{\omega}P_{\Lambda}$
on $\ell^{2}(\Lambda)$. We are interested in the matrix elements of its resolvent $G_{\omega,N}^{E}=({\mathcal{H}}_{\omega,N}-\tilde{E})^{-1}$ which are denoted by $G_{\omega,N}^{E}(j,k)$.
Let us denote the standard basis in $\ell^{2}({\mathbb{Z}})$ by $\{e_{i}\}$ and set $u=G^{E}_{\omega,N}e_{k}$. Assume that $0\leqslant j\leqslant k\leqslant N-1$. Then
$\begin{split}u(j)\,\,{\det({\mathcal{H}}_{\omega,N}-\tilde{E})}=({\mathcal{H}}_{\omega,N}-\tilde{E})e_{0}\wedge\dots{\\ }\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{j-1}\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})u\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{j+1}\wedge\dots({\mathcal{H}}_{\omega,N}-\tilde{E})e_{N-1}=\\ ({\mathcal{H}}_{\omega,N}-\tilde{E})e_{0}\wedge\dots\wedge(-\tilde{E}){\mathcal{H}}_{\omega,N}e_{j-1}\wedge e_{k}\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{j+1}\wedge\dots({\mathcal{H}}_{\omega,N}-\tilde{E})e_{N-1}=\\ (-1)^{j-k+1}({\mathcal{H}}_{\omega,N}-\tilde{E})e_{0}\wedge\dots\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{j-1}\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{j+1}\wedge\dots{\\ }\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{k}\wedge e_{k}\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{k+1}\dots\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{N-1}\\ =(-1)^{j-k+1}({\mathcal{H}}_{\omega,N}-\tilde{E})e_{0}\wedge\dots\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{j-1}\wedge\alpha_{j}e_{j}\wedge\dots{\\ }\wedge\alpha_{k-1}e_{k-1}\wedge e_{k}\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{k+1}\dots\wedge({\mathcal{H}}_{\omega,N}-\tilde{E})e_{N-1}\end{split}$
where $\alpha_{k}=\frac{2\omega_{k}}{\sqrt{(\omega_{k-1}+\omega_{k})(\omega_{k}+\omega_{k+1})}}$. One should mention that we have the agreement that $\det({\mathcal{H}}_{\omega,0}-\tilde{E})=1$. Consequently,
$G_{\omega,N}^{E}(j,k)=\frac{(-1)^{k-j+1}\det({\mathcal{H}}_{\omega,j}-\tilde{E})\det({\mathcal{H}}_{T^{k+1}\omega,N-k-1}-\tilde{E})}{\det({\mathcal{H}}_{\omega,N}-\tilde{E})}\prod_{i=j}^{k-1}\alpha_{i},$ (3.14)
for $0\leqslant j\leqslant k\leqslant N-1$. Note that
$\prod_{i=j}^{k-1}\alpha_{i}\leqslant\frac{2\sqrt{\omega_{j}\omega_{k-1}}}{\sqrt{(\omega_{j-1}+\omega_{j})(\omega_{k}+\omega_{k-1})}}\leqslant 2.$
Another important relation is the one that expresses $A^{E}_{N}(\omega)$ in terms of determinants considered above. Let $u$ be the formal solution of the equation $H_{\omega}u=Eu$ satisfying the intitial conditions u(1)=0u(-1)=0 and $u(0)=1$. Define $w(n)=\sqrt{\omega_{n}+\omega_{n-1}}\,u(n)$. Then
$({\mathcal{H}}_{\omega,N}-\tilde{E})w=-\alpha_{N-1}w(N)e_{N-1}\implies\sqrt{\frac{\omega_{0}+\omega_{-1}}{\omega_{N}+\omega_{N-1}}}=-\alpha_{N-1}u(N)G^{E}_{\omega,N}(0,N-1).$
Therefore,
$u(N)=\sqrt{\frac{\omega_{0}+\omega_{-1}}{\omega_{N}+\omega_{N-1}}}\det(\tilde{E}-{\mathcal{H}}_{\omega,N})\prod_{i=0}^{N-1}\alpha_{i}^{-1}.$
Let now $v$ be the solution satisfying the intitial conditions v(1)=1v(-1)=1 and $v(0)=0$. Then $\omega_{0}v(1)/\omega_{-1}=1$. Therefore,
$\frac{\omega_{0}}{\omega_{-1}}v(N)=\sqrt{\frac{\omega_{1}+\omega_{0}}{\omega_{N}+\omega_{N-1}}}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-1})\prod_{i=1}^{N-1}\alpha_{i}^{-1}.$
The representations of $u$ and $v$ lead to
$A^{E}_{N}(\omega)=\sqrt{\frac{\omega_{N-1}}{\omega_{-1}}}\begin{pmatrix}\frac{1}{\sqrt{\omega_{N}+\omega_{N-1}}}&0\\ &\\ 0&\frac{1}{\sqrt{\omega_{N-1}+\omega_{N-2}}}\end{pmatrix}\,\tilde{A}^{E}_{N}(\omega)\begin{pmatrix}\sqrt{\omega_{0}+\omega_{-1}}&0\\ &\\ 0&\sqrt{\omega_{-1}+\omega_{-2}}\end{pmatrix}$, (3.15)
where
$\tilde{A}^{E}_{N}(\omega)=\begin{pmatrix}\det(\tilde{E}-{\mathcal{H}}_{\omega,N})\prod_{i=0}^{N-1}\alpha_{i}^{-1}&-\alpha_{-1}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-1})\prod_{i=0}^{N-1}\alpha_{i}^{-1}\\ \,&\,\\ \det(\tilde{E}-{\mathcal{H}}_{\omega,N-1})\prod_{i=0}^{N-2}\alpha_{i}^{-1}&-\alpha_{-1}\det(\tilde{E}-{\mathcal{H}}_{T\omega,N-2})\prod_{i=0}^{N-2}\alpha_{i}^{-1}\end{pmatrix}$.
Combining (3.14) and (3.15) , we obtain
$|G_{\omega,N}^{E}(j,k)|\leqslant C_{\ell}\frac{\|A^{E}_{j}(\omega)\|\|A^{E}_{N-k}(T^{k}\omega)\|}{|\det({\mathcal{H}}_{\omega,N}-\tilde{E})|}\prod_{i=0}^{N-1}\alpha_{i},$ (3.16)
for $0\leqslant j\leqslant k\leqslant N-1$ with some constant $C_{\ell}$ that depends only on $\ell$.
Corollary 3.6.
For any $\varepsilon\in(0,1)$ and $\omega\in\Omega(\varepsilon)$ there is an $n_{1}=n_{1}(\omega,\varepsilon)$ such that the following is true:
1) For all $E\in I$,
$\frac{1}{n}\ln\|A^{E}_{n}(T^{s}\omega)\|\leqslant L(E)+2\varepsilon$ (3.17)
for all $n\in{\mathbb{N}}$ and $s_{0}\in{\mathbb{Z}}$ satisfying $n\geqslant\max\{n_{1},\ln^{10/7}(|s_{0}|+1)\}$.
2) For all $n\in{\mathbb{N}}$ and $s_{0}\in{\mathbb{Z}}$ satisfying $n\geqslant\varepsilon^{-1}\max\{n_{1},8\ln^{10/7}(|s_{0}|+1)\}$,
$|G^{E}_{T^{s_{0}}\omega,n}(j,k)|\leqslant C_{\ell}\frac{\exp((n-|j-k|)L(E)+C_{0}\varepsilon n)}{|\det({\mathcal{H}}_{T^{s_{0}}\omega,n}-\tilde{E})|}$ (3.18)
provided $E\notin\sigma({\mathcal{H}}_{\omega,n})$ and $j,k\in[0,n-1]\cap{\mathbb{Z}}$. The constant $C_{0}$ in this inequality depends only on $\Gamma$, while the constant $C_{\ell}$ depends only on $\ell$.
Proof. Set $m=\lceil n^{1/5}\rceil$. Then
$A^{E}_{n}(T^{s_{0}}\omega)=[A^{E}(T^{s_{0}+n}\omega)]^{-1}\cdots[A^{E}(T^{s_{0}+m^{5}-1}\omega)]^{-1}\prod_{s=0}^{m^{4}-1}A^{E}_{m}(T^{s_{0}+sm}\omega).$
Therefore , since the number of the factors containing inverse operators is $m^{5}-n\leqslant 31m^{4}$, we obtain
$\|A^{E}_{n}(T^{s_{0}}\omega)\|\leqslant\Gamma^{31m^{4}}\prod_{s=0}^{m^{4}-1}\|A^{E}_{m}(T^{s_{0}+sm}\omega)\|.$ (3.19)
Choose $n_{1}$ so that
$n_{1}\geqslant\max\{(\tilde{n}_{0}(\omega,\varepsilon))^{5},18\varepsilon^{-1},93^{5}\}\quad\text{and}\quad\frac{124\ln\Gamma}{n_{1}^{1/5}-31}<\varepsilon,$ (3.20)
where $\tilde{n}_{0}$ is the same as in Proposition 3.5. Now we apply Proposition 3.5 with $m$ replacing $n$ in it. Observe that for any $n$ satisfying the conditions of the first part of the corollary, $m\geqslant\max\{\tilde{n}_{0}(\omega,\varepsilon),\,\ln^{2/7}(|s_{0}|+1)\}$. Therefore,
$\begin{split}\frac{1}{n}\ln\|A^{E}_{n}(T^{s_{0}}\omega)\|\leqslant\frac{31m^{4}\ln\Gamma}{n}+\frac{m^{5}}{n}(L(E)+\varepsilon)\\ \leqslant\frac{62\ln\Gamma}{m-31}+\frac{m\varepsilon}{m-31}+L(E)\\ \leqslant L(E)+2\varepsilon.\end{split}$ (3.21)
Thus, relation (3.17) is established.
To prove (3.18), we need to estimate the product
$\|A^{E}_{j}(T^{s_{0}}\omega)\|\cdot\|A^{E}_{n-k}(T^{s_{0}+k}\omega)\|.$
Suppose $n>\varepsilon^{-1}\max\{n_{1},8\ln^{10/7}(|s_{0}|+1)\}$. Set $h=\lceil\varepsilon n\rceil$. Then for $j\geqslant 0$, we have
$\|A^{E}_{j}(T^{s_{0}}\omega)\|\leqslant\|A^{E}_{j+h}(T^{s_{0}-h}\omega)\|\cdot\|[A^{E}_{h}(T^{s_{0}-h}\omega)]^{-1}\|.$
Observe that $j+h\geqslant h\geqslant n_{1}$. Therefore, to apply (3.17), we only need to check that $h>\ln^{10/7}(|s_{0}-h|+1)$, which can be easily derived from the inequalities
$\begin{split}\ln^{10/7}(|a|+|b|+1)\leqslant 2^{10/7}(\ln^{10/7}(|a|+1)+\ln^{10/7}(|b|+1)),\\ \text{and}\qquad 2^{10/7}\ln^{10/7}(b+1)\leqslant b,\qquad\text{for }\quad b>10.\end{split}$
Thus,
$\|A^{E}_{j}(T^{s_{0}}\omega)\|\leqslant\exp((j+2h)(L(E)+2\varepsilon))\leqslant\exp(L(E)j+C_{0}\varepsilon n),$ (3.22)
with the constant $C_{0}$ depending only on $\Gamma$.
Similarly,
$\|A^{E}_{n-k}(T^{s_{0}+k}\omega)\|\leqslant\|A^{E}_{n-k+h}(T^{s_{0}+k-h}\omega)\|\cdot\|[A^{E}_{h}(T^{s_{0}+k-h}\omega)]^{-1}\|$
This time we need to check that $h\geqslant\ln^{10/7}(|s_{0}+k-h|+1)$. For that purpose, it is sufficient to note that
$2^{10/7}\ln^{10/7}(|-k+h|+1)\leqslant 2^{10/7}\ln^{10/7}(n+1)\leqslant\frac{\varepsilon n}{2}\leqslant\frac{h}{2}$
is implied by the condition $n\geqslant 18\varepsilon^{-1}$. Consequently,
$\|A^{E}_{n-k}(T^{s_{0}+k}\omega)\|\leqslant\exp((n-k+2h)(L+2\varepsilon))\leqslant\exp((n-k)L(E)+C_{0}\varepsilon n)$ (3.23)
Multiplying (3.22) by (3.23), we obtain (3.18) for $j\leqslant k$. Taking the adjoint operator, we obtain it for $j\geqslant k$ as well. $\,\,\,\Box$
In [2], Avila, Damanik, and Zhang introduce a process termed the ”elimination of double resonances.” We adopt a specific, context-dependent usage of ”resonances,” referring to points $\omega$ rather than energy levels. It can be shown that this set of resonant points in $\Omega$ has small measure.
For $N\in{\mathbb{Z}}_{+}$, we define
$\bar{N}=[N^{\ln N}].$
For a given $\varepsilon$ and $N\in{\mathbb{Z}}_{+}$, the set $D_{N}=D_{N}(\varepsilon)$ is defined as the set of $\omega\in\Omega$ such that
$\|G^{E}_{T^{s}\omega,[-N_{1},N_{2}]}\|\geqslant e^{K^{2}}$ (3.24)
and
$\frac{1}{m}\ln\|A_{m}^{E}(T^{s+r}\omega)\|\leqslant L(E)-\varepsilon$ (3.25)
for some choice of $s\in{\mathbb{Z}}$, $K\geqslant\max\{N,\ln^{2}(|s|+1)\}$, $E\in I$, $0\leqslant N_{1},N_{2}\leqslant K^{9}$, $K^{10}\leqslant r\leqslant\bar{K}$, and $m\in\{K,2K\}$ (it is either $K$ or $2K$).
Proposition 3.7.
For any $\varepsilon>0$ there are positive constants $C,\eta>0$ for which
$\mu(D_{N}(\varepsilon))\leqslant Ce^{-\eta N},\qquad\forall N\in{\mathbb{Z}}_{+}.$
To prove Proposition 3.7, we need the following lemma from [2]. For an admissible $\underline{l}=(l_{1},\dots,l_{n})$, define $\Omega^{+}_{\underline{l}}=[0;l]\cap\Omega^{+}$. Then we define
[…]
Lemma 3.8.
Assume that $\mu$ has a bounded distortion property. Then there is a constant $C\geqslant 1$ such that
$\frac{d\mu^{+}_{\overline{l}}}{d\mu^{+}}\leqslant C\qquad\text{a.e. on}\quad\Omega^{+},$ (3.27)
uniformly in all admissible $\overline{l}$. In particular,
$\int_{\Omega_{+}}fd\mu^{+}_{\overline{l}}\leqslant C\int_{\Omega_{+}}fd\mu^{+}.$ (3.28)
Proof of Proposition 3.7. The proof follows a similar logic to that in [2], with the modification that we occasionally replace the operator $H_{\omega}$ by ${\mathcal{H}}_{\omega}$ and the point $E$ by $\tilde{E}=2\cos\sqrt{E}$.
Let $N$ and $\varepsilon$ be given. Fix $s\in{\mathbb{Z}}$ and $K\geqslant\max\{N,\ln^{2}(|s|+1)\}$. Define the set $D_{K,s}$ as the set of $\omega$ for which (3.24),(3.25) are satisfied for some $N_{1}$, $N_{2}$, $r$, and $m$ obeying the conditions $0\leqslant N_{1},N_{2}\leqslant K^{9}$, $K^{10}\leqslant r\leqslant\bar{K}$, and $m\in\{K,2K\}$. Note that
$D_{K,s}=\bigcup_{0\leqslant N_{1},N_{2}\leqslant K^{9}}\bigcup_{K^{10}\leqslant r\leqslant\bar{K}}D_{1}(N_{1},N_{2},r,s)\cup D_{2}(N_{1},N_{2},r,s)$ (3.29)
where $D_{j}(N_{1},N_{2},r,s)$ is the set of $\omega$ for which there is an $E\in I$ such that (3.24),(3.25) hold for $m=jK$.
To estimate the measure $\mu(D_{1})$, let us assume that $\omega\in D_{1}(N_{1},N_{2},r,s)$. That means (3.24) and (3.25) hold for some $E\in I$. Consequently, there is an $\tilde{E}_{0}\in\sigma({\mathcal{H}}_{T^{s}\omega,[-N_{1},N_{2}]})$ for which
$|\tilde{E}-\tilde{E}_{0}|\leqslant\|G^{E}_{T^{s}\omega,[-N_{1},N_{2}]}\|^{-1}\leqslant e^{-K^{2}},\quad\text{with}\quad\tilde{E}=2\cos\sqrt{E}.$ (3.30)
Denote also by $E_{0}$ the unique sulution of the equation $\tilde{E}_{0}=2\cos\sqrt{E}_{0}$ on $I_{0}$ where $I_{0}$ is the unique interval of the form $(\pi n)^{2},(\pi(n+1))^{2})$ containing $I$. Using the inequality
$|g_{n}(\omega,E)-g_{n}(\omega,E_{0})|\leqslant C\Gamma^{n-1}|\tilde{E}-\tilde{E}_{0}|$
and selecting $K$ (i.e. $N$) so large that $C\Gamma^{K}e^{-K^{2}}<\varepsilon/6$, we conclude that
$g_{K}(T^{s+r}\omega,E_{0})\leqslant g_{K}(T^{s+r}\omega,E)+\frac{\varepsilon}{6}\leqslant L(E)-\frac{5\varepsilon}{6}$
Applying Hölder continuity of $L(E)$ stated in (3.9) and selecting $K$ so that $Ce^{-\beta K^{2}}<\varepsilon/6$, we derive the estimate
$g_{K}(T^{s+r}\omega,E_{0})\leqslant L(E_{0})-\frac{2\varepsilon}{3}.$ (3.31)
Thus, we see that
$D_{1}(N_{1},N_{2},r,s)\subset\hat{D}(N_{1},N_{2},r,s)$ (3.32)
where $\hat{D}(N_{1},N_{2},r,s)$ is the set of $\omega$ for which
$g_{K}(T^{s+r}\omega,E_{0})\leqslant L(E_{0})-\frac{2\varepsilon}{3}$
for some $E_{0}\in I_{0}$ such that $2\cos\sqrt{E_{0}}\in\sigma({\mathcal{H}}_{T^{s}\omega,[-N_{1},N_{2}]})$. The measure of the set $\hat{D}$ is the same as the measure of the set
$T^{s}(\hat{D})=\bigcup_{\tilde{E}_{0}\in\sigma({\mathcal{H}}_{\omega,[-N_{1},N_{2}]})}\{\omega:\,\,g_{K}(T^{r}\omega,E_{0})\leqslant L(E_{0})-\frac{2\varepsilon}{3}\}$ (3.33)
where $\tilde{E}_{0}=2\cos E_{0}$. Let $\underline{l}$ be an admissible multiindex such that $|\underline{l}|=K^{2}+1$. Define $\Omega_{\underline{l}}$ to be the set $\Omega_{\underline{l}}=[-K^{2};\underline{l}]$. For each such $\underline{l}$, we fix $\omega^{(\underline{l})}\in\Omega_{\underline{l}}$. Then all points $\omega\in\Omega_{\underline{l}}$ satisfy the condition $d(\omega,\omega^{(\underline{l})})\leqslant e^{-K^{2}}$. Therefore, for each $\tilde{E}_{0}\in\sigma({\mathcal{H}}_{\omega,[-N_{1},N_{2}]})$ there is an $\tilde{E}^{\prime}\in\sigma({\mathcal{H}}_{\omega^{(\underline{l})},[-N_{1},N_{2}]})$ such that
$|\tilde{E}_{0}-\tilde{E}^{\prime}_{0}|\leqslant\|{\mathcal{H}}_{\omega^{(\underline{l})},[-N_{1},N_{2}]}-{\mathcal{H}}_{\omega,[-N_{1},N_{2}]}\|\leqslant Ce^{-\alpha K^{2}}$ (3.34)
Repeating the arguments that lead us to (3.31), we obtain that for $\omega\in\Omega_{\underline{l}}$,
$g_{K}(T^{r}\omega,E_{0})<L(E_{0})-\frac{2\varepsilon}{3}\implies g_{K}(T^{r}\omega,E^{\prime})<L(E^{\prime})-\frac{\varepsilon}{3}$ (3.35)
for some $\tilde{E}^{\prime}\in\sigma({\mathcal{H}}_{\omega^{(\underline{l})},[-N_{1},N_{2}]})$ provided $K$ satisfies the inequalities $Ce^{-\alpha\beta K^{2}}<\varepsilon/6$ and $C\Gamma^{K}e^{-\alpha K^{2}}<\varepsilon/6$. Here, as before, $\tilde{E}^{\prime}=2\cos\sqrt{E^{\prime}}$, and $E^{\prime}\in I_{0}$.
Put differently,
$\hat{D}\subset\bigcup_{\underline{l}}\,\,\bigcup_{\tilde{E}^{\prime}\in\sigma({\mathcal{H}}_{\omega^{(\underline{l})},[-N_{1},N_{2}]})}\{\omega\in\Omega_{\underline{l}}:\,\,g_{K}(T^{r}\omega,E^{\prime})<L(E^{\prime})-\frac{\varepsilon}{3}\}.$ (3.36)
Define now $S_{K}(E,\varepsilon)$ to be the set $\{\omega:\,\,g_{K}(\omega,E)<L(E)-\varepsilon\}$. Then the set in the braces on the right hand side of (3.36) is the intersection
$\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[S_{K}(E^{\prime},\varepsilon/3)].$
To estimate the measure of this set, we introduce
$\tilde{S}(E)=\bigcup_{\omega\in T^{-K^{2}}S_{K}(E,\varepsilon/3)}W_{\rm loc}^{s}(\omega)$
Note that if $\omega^{\prime}\in W^{s}_{\rm loc}(\omega)$, then $d(T^{K^{2}}\omega,T^{K^{2}}\omega^{\prime})\leqslant e^{-K^{2}}$. Consequently, $g_{K}(T^{K^{2}}\omega,E)=g_{K}(T^{K^{2}}\omega^{\prime},E)$ for $K>1$, and therefore,
$g_{K}(T^{K^{2}}\omega^{\prime},E)\leqslant L(E)-\frac{\varepsilon}{3}.$
This implies that
$T^{-K^{2}}S_{K}(E,\varepsilon/3)=\tilde{S}(E).$
Clearly,
$T^{-K^{2}}\bigl(\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[S_{K}(E^{\prime},\varepsilon/3)]\bigr)=T^{-K^{2}}\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[\tilde{S}(E^{\prime})]$ (3.37)
which is locally $s$-saturated.
Denote now $S^{+}(E)=\pi^{+}(\tilde{S}(E))$ and observe that $\pi^{+}\bigl(T^{-K^{2}}\Omega_{\underline{l}}\bigr)=[0;\underline{l}]^{+}=\Omega_{\underline{l}}^{+}.$ Allss-locally saturated sets $X$ have the following two properties: $\mu(X)=\mu^{+}(\pi^{+}(X))$ and $\pi^{+}T^{-n}X=T_{+}^{-n}\pi^{+}X$. Therefore, since
$\pi^{+}\Bigl[T^{-K^{2}}\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[\tilde{S}(E^{\prime})]\Bigr]\subset\pi^{+}(T^{-K^{2}}\Omega_{\underline{l}})\,\,\bigcap\,\pi^{+}(T^{-r}[\tilde{S}(E^{\prime})]),$
we conclude that
$\begin{split}\mu\Bigl[T^{-K^{2}}\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[\tilde{S}(E^{\prime})]\Bigr]\leqslant\mu^{+}\Bigl(\pi^{+}(T^{-K^{2}}\Omega_{\underline{l}})\,\,\bigcap\,\pi^{+}(T^{-r}[\tilde{S}(E^{\prime})])\Bigr)=\\ \mu^{+}\Bigl(\Omega_{\underline{l}}^{+}\,\,\bigcap\,T_{+}^{-r}[S^{+}(E^{\prime})]\Bigr)\end{split}$
Thus, by $T$-invariance of the measure $\mu$, and the fact that $r>|\underline{l}|=K^{2}+1$,
$\begin{split}\mu\bigl(\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[S_{K}(E^{\prime},\varepsilon/3)]\bigr)\leqslant\mu^{+}\Bigl(\Omega_{\underline{l}}^{+}\,\,\bigcap\,T_{+}^{-r}[S^{+}(E^{\prime})]\Bigr)=\\ \mu^{+}(\Omega_{\underline{l}}^{+})\mu^{+}_{\underline{l}}(T_{+}^{-r+|\underline{l}|}[S^{+}(E^{\prime})])\leqslant C\mu(\Omega_{\underline{l}})\mu^{+}(T_{+}^{-r+|\underline{l}|}[S^{+}(E^{\prime})])=\\ C\mu(\Omega_{\underline{l}})\mu^{+}(S^{+}(E^{\prime}))=C\mu(\Omega_{\underline{l}})\mu(\tilde{S}(E^{\prime}))=C\mu(\Omega_{\underline{l}})\mu(S_{K}(E^{\prime},\varepsilon/3))\end{split}$
Applying the ULD property (on a slightly bigger interval than $I$), we obtain
$\mu\bigl(\Omega_{\underline{l}}\,\,\bigcap\,T^{-r}[S_{K}(E^{\prime},\varepsilon/3)]\bigr)\leqslant C_{0}\,\mu(\Omega_{\underline{l}})e^{-c\varepsilon K}.$
Since the number of the eigenvalues $\tilde{E}^{\prime}\in\sigma({\mathcal{H}}_{\omega^{(\underline{l})},[-N_{1},N_{2}]})$ does not exceed $2K^{9}+1$, we conclude that
$\mu(D_{1}(N_{1},N_{2},r,s))\leqslant\mu(\hat{D}(N_{1},N_{2},r,s))\leqslant C(2K^{9}+1)e^{-c\varepsilon K}\leqslant C_{\varepsilon}e^{-c_{\varepsilon}K}.$
Similarly, one can show that
$\mu(D_{2}(N_{1},N_{2},r,s))\leqslant C_{\varepsilon}e^{-c_{\varepsilon}K}.$
Summing up over all $N_{1}$, $N_{2}$ and $r$ obeying the conditions in the definition of $D_{N}(\varepsilon)$, we obtain that
$\mu(D_{K,s})\leqslant K^{18}\bar{K}C^{2}_{\varepsilon}e^{-2c_{\varepsilon}K}\leqslant\tilde{C}e^{-\tilde{\eta}K}.$
Finally, we use the relation
$D_{N}(\varepsilon)=\bigcup_{s\in{\mathbb{Z}}}\bigcup_{K\geqslant\max\{N,\ln^{2}(|s|+1)\}}D_{K,s}\subset\bigcup_{K\geqslant N}\bigcup_{|s|\leqslant e^{\sqrt{K}}}D_{K,s}$
to conclude that
$\mu(D_{N}(\varepsilon))\leqslant Ce^{-\eta N}.$
$\Box$
By Proposition 3.7, the set
$\Omega_{+}(\varepsilon)=\Omega\setminus\bigcap_{k=1}^{\infty}\bigcup_{N>k}(D_{N}(\varepsilon))$ (3.38)
is a set of full measure. The same is true about the set $\Omega(\varepsilon)$ from Proposition 3.5. Since the set $\Omega_{+}(\varepsilon)$ is decreasing in $\varepsilon$ and the set $\Omega(\varepsilon)$ is increasing in $\varepsilon$, the set
$\bigl(\bigcap_{\varepsilon\in(0,1)}\Omega_{+}(\varepsilon)\bigr)\bigcap\bigl(\bigcap_{\varepsilon\in(0,1)}\Omega(\varepsilon)\bigr)$
has full measure. Let us consider the intersection
$\Omega_{*}=\Omega_{I}\bigcap\bigl(\bigcap_{\varepsilon\in(0,1)}\Omega_{+}(\varepsilon)\bigr)\bigcap\bigl(\bigcap_{\varepsilon\in(0,1)}\Omega(\varepsilon)\bigr),$ (3.39)
where $\Omega_{I}$ is the set of all $\omega$ for which the spectrum of $H_{\omega}$ covers the interval $I$.
Let $\varepsilon\in(0,1)$, $\omega\in\Omega_{*}$ be given and let $u$ be a generalized eigenfunction of $H_{\omega}$ corresponding to $E\in\sigma(H_{\omega})$. We will assume that $u(s_{0})=1$ for some choice of $s_{0}\in\{0,1\}$. Define
$K=\lceil\frac{1}{\varepsilon}\max\{N,2\ln^{2}(|s_{0}|+1)\}\rceil$ (3.40)
where $N\geqslant N_{0}=\max\{\tilde{n}_{0},n_{1},n_{2}\}$ where $\tilde{n}_{0}$ and $n_{1}$ are taken from Proposition 3.5 and Corollary 3.6, while $n_{2}=n_{2}(\omega,\varepsilon)$ is the number such that $\omega\notin D_{N}(\varepsilon)$ for all $N\geqslant n_{2}$. Since $s_{0}\in\{0,1\}$, we can replace $\ln(|s_{0}|+1)$ by $\ln 2$ and assume that $K\sim N/\varepsilon$.
We consider two claims that hold for $\omega\in\Omega_{*}$ and are similar to the those in [2].
Claim 1. There are integer numbers $a_{1}$, $a_{2}$, $b_{1}$ and $b_{2}$ having the properties
$K^{9}\leqslant a_{1}\leqslant-K^{3}+1,\qquad 0\leqslant a_{2}\leqslant K^{9}$ (3.41)
and $b_{i}\in\{a_{i}+K^{3}-2,a_{i}+K^{3}-1,a_{i}+K^{3}\}$ such that
$|G^{E}_{T^{s_{0}}\omega,\Lambda_{i}}(j,k)|\leqslant\exp(-|j-k|L(E)+C_{0}\varepsilon K^{3})$ (3.42)
for all $j,k\in\Lambda_{i}=[a_{i},b_{i})\cap{\mathbb{Z}}$ where $C_{0}$ depends only on $\Gamma$.
Proof. We apply Proposition 3.5 twice with: $s_{0}$ and $s_{0}-K^{9}$. The parameter $n$ is set equal to $K^{3}$. As a result, we obtain
$L(E)-\frac{1}{K^{6}}\sum_{s=0}^{K^{6}-1}g_{K^{3}}(T^{s_{0}+sK^{3}}\omega,E)<\varepsilon$
and
$L(E)-\frac{1}{K^{6}}\sum_{s=0}^{K^{6}-1}g_{K^{3}}(T^{s_{0}+sK^{3}-K^{9}}\omega,E)<\varepsilon.$
Consequently, there are integers $s_{1}$ and $s_{2}$ in the intervals $K^{9}\leqslant s_{1}\leqslant-K^{3}$ and $0\leqslant s_{2}\leqslant K^{9}-K^{3}$ for which
$\frac{1}{K^{3}}\ln\|A^{E}_{K^{3}}(T^{s_{0}+s_{i}}\omega)\|>L(E)-\varepsilon.$
This leads to the estimate
$\frac{1}{K^{3}}(\ln 4+\ln|\det({\mathcal{H}}_{T^{s_{0}+a_{i}}\omega,k_{i}^{3}}-\tilde{E})|)>L(E)-\varepsilon$
for $\tilde{E}=2\cos\sqrt{E}$ and some choice of $k_{i}\in\{K^{3},K^{3}-1,K^{3}-2\}$ and $a_{i}\in\{s_{i},s_{i}+1\}$. Define $b_{i}=a_{i}+k_{i}$ and choose $N$ so large that $\ln 4/N_{0}^{3}<\varepsilon$. Then the preceding inequality will turn into the estimate
$\frac{1}{K^{3}}\ln|\det({\mathcal{H}}_{T^{s_{0}}\omega,\Lambda_{i}}-\tilde{E})|>L(E)-2\varepsilon,$
where $\Lambda_{i}=[a_{i},b_{i})$. Thus, we finially obtain that
$\begin{split}|G^{E}_{T^{s_{0}}\omega,\Lambda_{i}}(j,k)|\leqslant\frac{\exp(K^{3}-|j-k|L(E)+C_{0}\varepsilon K^{3})}{|\det({\mathcal{H}}_{T^{s_{0}}\omega,\Lambda_{i}}-\tilde{E})|}\leqslant\\ \exp(-|j-k|L(E)+(C_{0}+2)\varepsilon K^{3}),\end{split}$
for $j,k\in\Lambda_{i}$. To apply Corollary 3.6, we might need to enlarge $N_{0}$ so that $K^{3}>\frac{4}{\varepsilon}\ln^{10/7}(K^{9}+2)$ for all $K>N_{0}$ $\,\Box$
Claim 2. Let $2\varepsilon C_{0}<L(E)$. Let $u$ be the generalized solution of the equation $H_{\omega}u=Eu$ satisfying the conditions $|u(n)|\leqslant C_{u}(1+|n|)$ for all $n\in{\mathbb{Z}}$ and $u(s_{0})=1$. Let $\Lambda_{i}$ be the same as in Claim 1. Define
$\ell_{i}=\Bigl[\frac{a_{i}+b_{i}}{2}\Bigr].$
Then
$|u(s_{0}+\ell_{i})|\leqslant e^{-2K^{2}},\qquad i=1,2.$ (3.43)
whenever $N>N_{0}$ where $N_{0}$ is sufficiently large. Here, $N_{0}$ depends on $\omega$, $\varepsilon$ , and $C_{u}$.
Proof. Again, we define $w(n)=\sqrt{\omega_{n}+\omega_{n-1}}u(n)$. Then for any interval $[a,b]$,
$w(n)=-G_{\omega,[a,b]}^{E}(n,a)\alpha_{a-1}w(a-1)-G^{E}_{\omega,[a,b]}(n,b)\alpha_{b}w(b+1)$ (3.44)
for all $a\leqslant n\leqslant b$. In particular, we may apply this relation for $n=\ell_{i}+s_{0}$ with $\ell_{i}=[(a_{i}+b_{i})/2]$ and the interval $\Lambda_{i}+s_{0}$. As a result, we obtain
$\begin{split}|w(\ell_{i}+s_{0})|\leqslant 2(|w(a_{i}+s_{0})|+|w(b_{i}+s_{0}-1)|)\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3})\\ \leqslant C_{u}(5+|a_{i}|+|b_{i}|)\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3})\\ \leqslant C_{u}(5+2K^{9})\exp(-(K^{3}/2-1)L(E)+C_{0}\varepsilon K^{3}).\end{split}$
Thus, (3.43) holds for large $K$ and small $\varepsilon$ $\,\Box$
Now we use the condition $|u(s_{0})|=1$ and apply (3.44) with $a=s_{0}+\ell_{1}+1$ and $b=s_{0}+\ell_{2}-1$ to get the estimate
$\begin{split}1\leqslant|w(s_{0})|\leqslant 2|G^{E}_{\omega,[a,b]}(s_{0},a)|\,|w(s_{0}+\ell_{1})|+2|G^{E}_{\omega,[a,b]}(s_{0},b)|\,|w(s_{0}+\ell_{2})|\\ \leqslant 2\bigl(|G^{E}_{\omega,[a,b]}(s_{0},a)|+|G^{E}_{\omega,[a,b]}(s_{0},b)|\bigr)e^{-2K^{2}}\leqslant 4\|G^{E}_{\omega,[a,b]}\|e^{-2K^{2}}.\end{split}$
Consequently,
$\|G^{E}_{\omega,[a,b]}\|\geqslant e^{K^{2}}.$
This condition (3.24) holds, because $0\leqslant-a,b\leqslant K^{9}$. Since $\omega\notin D_{N}(\varepsilon)$, we conclude that (3.25) fails. That is,
$\frac{1}{m}\ln\|A^{E}(T^{s_{0}+r}\omega)\|\geqslant L(E)-\varepsilon,$ (3.45)
for all $N>N_{0}$, $K^{10}\leqslant r\leqslant\bar{K}$, and $m\in\{K,2K\}$.
To proceed, we introduce the following lemma, known as the Avalanche Principle, developed by Goldstein and Schlag to study the Lyapunov exponents of Schrödinger cocycles.
Lemma 3.9.
(Avalanche Principle). Let $A^{(1)},\dots,A^{(n)}$ be ${\rm SL}(2,{\mathbb{R}})$ matrices satisfying the following conditions:
$\|A(j)\|\geqslant\lambda>n,\qquad\forall 1\leqslant j\leqslant n$
and
$\bigl|\ln\|A^{(j+1)}\|+\ln\|A^{(j)}\|-\ln\|A^{(j+1)}A^{(j)}\|\bigr|<\frac{1}{2}\ln\lambda,\qquad\forall 1\leqslant j\leqslant n-1.$
Then
$\Bigl|\ln\|A^{(n)}\dots A^{(1)}\|+\sum_{j=2}^{n-1}\ln\|A^{(j)}\|-\sum_{j=1}^{n-1}\ln\|A^{(j+1)}A^{(j)}\|\leqslant C\frac{n}{\lambda}.$ (3.46)
Define now
$A^{(j)}=A^{E}_{K}(T^{s_{0}+K^{10}+(j-1)K}\omega)$ (3.47)
for $1\leqslant j\leqslant n$ where $K^{10}\leqslant n\leqslant K^{-1}\bar{K}-K^{9}$. Then according to (3.45),
$\|A(j)\|\geqslant\lambda=\exp(K(L(E)-\varepsilon)).$ (3.48)
Moreover, $\lambda>n$ if $N_{0}$ is sufficiently large. Since $\omega\in\Omega(\varepsilon)$ from Corollary 3.6, we have
$\frac{1}{n}\ln\|A_{n}^{E}(T^{s}\omega)\|\leqslant L(E)+2\varepsilon$
for all $n>\max\{n_{1},\ln^{10/7}(|s_{0}|+1)\}$ which implies that
$\|A^{(j)}\|\leqslant\exp(K(L(E)+2\varepsilon))$
as long as $K>\ln^{10/7}(\bar{K}+1)$. Thus, (3.45) with $m=2K$ implies
$\begin{split}0\leqslant\ln\|A^{(j+1)}\|+\ln\|A^{(j)}\|-\ln\|A^{(j+1)}A^{(j)}\|\leqslant\\ 2K(L(E)+2\varepsilon)-2K(L(E)-2\varepsilon)=6K\varepsilon\leqslant\frac{1}{2}\ln\lambda=\frac{K(L(E)+2\varepsilon)}{2},\end{split}$ (3.49)
if $10\varepsilon<L(E)$.
So all conditions of Lemma 3.9 are satisfied. Denoting $\hat{N}=nK$ and $r_{0}=K^{10}$, we write the resulting inequality in the form
$\begin{split}\ln\|A^{E}_{\hat{N}}(T^{s_{0}+r_{0}}\omega)\|=\ln(\|A^{(n)}\dots A^{(1)}\|)\geqslant\sum_{j=1}^{n-1}\ln\|A^{(j+1)}A^{(j)}\|-\sum_{j=2}^{n-1}\ln\|A^{(j)}\|-C\\ \geqslant 2(n-1)(L(E)-2\varepsilon)K-(n-2)(L(E)+2\varepsilon)K-C\geqslant\\ \hat{N}(L(E)-5\varepsilon)\end{split}$
if $N_{0}$ is sufficiently large. Here $\hat{N}$ is any number of the form $nK$ between $K^{11}$ and $\bar{K}-K^{10}$.
Let us try to replace $\hat{N}$ by an arbitrary sufficiently large number $\ell$. Observe that the union of intervals $K^{11}+K^{10},\bar{K}]$ contains all sufficiently large integer numbers. Therefore, we may assume that $\ell\in[K^{11}+K^{10},\bar{K}]$ is of the form $\ell=nK+p$ where $0\leqslant p<K$ and $K^{10}+K^{9}\leqslant n\leqslant K^{-1}\bar{K}-1$. In this case,
$\begin{split}\|A^{E}_{\ell}(T^{s_{0}}\omega)\|\geqslant\frac{\|A^{E}_{\ell-r_{0}}(T^{s_{0}+r_{0}}\omega)\|}{\|[A^{E}_{r_{0}}(T^{s_{0}}\omega)]^{-1}\|}\geqslant\frac{\exp((\ell-r_{0})(L(E)-5\varepsilon))}{\Gamma^{r_{0}}}\\ \geqslant\exp(\ell(L(E)-6\varepsilon)),\end{split}$
for all sufficiently large $N_{0}$. This implies the relation
$\liminf_{n\to\infty}\frac{1}{n}\ln\|A^{E}_{n}(T^{s_{0}}\omega)\|\geqslant L(E)-6\varepsilon,$
for $s_{0}\in\{0,1\}$.
Let ${\mathfrak{G}}(H_{\omega})$ be the set of energies $E$ for which the operator $H_{\omega}$ has a generalized eigenfunction. We have established the following fact:
Theorem 3.10.
There is a subset of full measure $\Omega_{0}\subset\Omega$, such that for every $\omega\in\Omega_{0}$, and every $E\in{\mathfrak{G}}(H_{\omega})\cap I$,
$\lim_{n\to\infty}\frac{1}{n}\ln\|A^{E}_{n}(\omega)\|=\lim_{n\to-\infty}\frac{1}{|n|}\ln\|A^{E}_{n}(\omega)\|=L(E).$ (3.50)
It was shown in [10] that this statement implies Theorem 1.1.
4. Hölder continuity of the Lyapunov exponent
The main credit for the results of this section belongs to the authors of [10] as the core methodology was established in their original paper. We only need to adjust it to the specific setting.
Let us define $L_{n}(E)$ as the integral
$L_{n}(E)=\frac{1}{n}\int_{\Omega}\ln\|A_{n}^{E}(\omega)\|\,d\mu$ (4.1)
The following lemma is proved in [10] for a different cocycle, but one can easily see that the proof works in our case too.
Lemma 4.1.
There are constants $c_{0},C_{0}>0$ that depend on $\Gamma$ and the set $I$, such that
$|L(E)-L_{n}(E)-2L_{2n}(E)|\leqslant C_{0}e^{-c_{0}n},\qquad\text{for all}\quad E\in I.$ (4.2)
Proof. First, choose $\varepsilon>0$ so small that
$0<\frac{4\varepsilon}{\gamma-\varepsilon}<\frac{1}{2},\qquad\text{where}\quad\gamma=\sup_{E\in I}L(E).$
Without loss of generality, we may assume that the constant $c>0$ in LDE satisfies the condition
$c<4(\gamma-\varepsilon).$
Since we only need to prove (4.1) for large values of $n$, we may assume that $(\ln n)/c>1$. Then there is an integer number $N$ such that
$e^{cN/5}\leqslant n\leqslant e^{cN/4},\quad\text{which is smaller than}\quad e^{(L-\varepsilon)N}.$ (4.3)
We will also assume that $2nCe^{-cN}\leqslant e^{-cN/4}$. For each $\omega\in\Omega$, consider the matrices
$A^{(j)}(\omega)=A^{E}_{N}(T^{(j-1)N}\omega)$
According to the ULD, if $B(n)$ is the set of $\omega$’s on which one of the inequalities
$\Bigl|\frac{1}{N}\ln\|A^{(j)}(\omega)\|-L\Bigr|<\varepsilon,\qquad\Bigl|\frac{1}{N}\ln\|A^{(j+1)}(\omega)A^{(j)}(\omega)\|-L\Bigr|<\varepsilon$ (4.4)
does not hold for some $1\leqslant j\leqslant n$, then
$\mu(B(n))\leqslant e^{-cN/4}.$ (4.5)
In particular,
$\|A^{(j)}\|\geqslant e^{(L-\varepsilon)N}>n$
and
$\Bigl|\ln\|A^{(j+1)}(\omega)\|+\ln\|A^{(j)}(\omega)\|-\ln\|A^{(j+1)}(\omega)A^{(j)}(\omega)\|\Bigr|<4N\varepsilon$
for all $\omega\notin B(n)$. Applying the Avalanche Principle with $\lambda=e^{(L-\varepsilon)N}$, we obtain that
$\Bigl|\ln\|A^{E}_{nN}(\omega)\|+\sum_{j=2}^{n-1}\ln\|A_{N}^{E}(T^{(j-1)N})\|-\sum_{j=1}^{n-1}\ln\|A_{2N}^{E}(T^{(j-1)N})\|\Bigr|\leqslant C\frac{n}{\lambda}.$
Conseqently,
$\Bigl|L_{nN}+\frac{(n-2)}{n}L_{N}-\frac{2(n-1)}{n}L_{2N}\Bigr|\leqslant\frac{C}{\lambda N}+4\mu(B(n))\ln\Gamma\leqslant Ce^{-cN/4}.$
which leads to the inequality
$\Bigl|L_{nN}+L_{N}-2L_{2N}\Bigr|\leqslant Ce^{-cN/4}+\frac{4}{n}\ln\Gamma\leqslant Ce^{-cN/5}.$
Now, consider a sequence of indices $n_{s}$ such that $n_{s+1}$ is an integer multiple of $n_{s}$, and
$e^{cn_{s}/5}\leqslant\frac{n_{s+1}}{n_{s}}\leqslant\frac{1}{2}e^{cn_{s}/4}.$
Setting $N=n_{s}$ and $n$ equal to either $\frac{n_{s+1}}{n_{s}}$ or $2\frac{n_{s+1}}{n_{s}}$, we obtain two estimates
$\Bigl|L_{n_{s+1}}+L_{n_{s}}-2L_{n_{s}}\Bigr|\leqslant Ce^{-cn_{s}/5},$ (4.6)
and
$\Bigl|L_{2n_{s+1}}+L_{n_{s}}-2L_{n_{s}}\Bigr|\leqslant Ce^{-cn_{s}/5}.$ (4.7)
Combining these inequalities (4.6) and (4.7), we obtain
$\Bigl|L_{2n_{s+1}}-L_{n_{s+1}}\Bigr|\leqslant 2Ce^{-cn_{s}/5}.$ (4.8)
Now we see that (4.6) and (4.8) imply
$\Bigl|L_{n_{s+1}}-L_{n_{s}}\Bigr|\leqslant\tilde{C}e^{-cn_{s-1}/5},\qquad\text{for}\quad s\geqslant 2.$
Therefore,
$\Bigl|L-L_{n_{2}}\Bigr|=\Bigl|\sum_{s=2}^{\infty}(L_{n_{s+1}}-L_{n_{s}})\Bigr|\leqslant Ce^{-cn_{1}/5}$
Thus, writing (4.6) with $s=1$ and replacing $L_{n_{2}}$ by $L$ in the resulting inequality, we obtain
$\Bigl|L+L_{n_{1}}-2L_{n_{1}}\Bigr|\leqslant Ce^{-cn_{1}/5}.$
It remains to observe that the first member $n_{1}$ of the sequence of indices was an arbitrary sufficiently large number. $\,\Box$
Theorem 4.2.
There is a constant $C>0$ and an exponent $\beta>0$ such that
$|L(E)-L(E^{\prime})|\leqslant C|E-E^{\prime}|^{\beta}$ (4.9)
for all $E,E^{\prime}\in I$.
Proof. Observe that
$|L_{n}(E)-L_{n}(E^{\prime})|\leqslant C_{\ell,I}\Gamma^{n-1}|E-E^{\prime}|$
Thus, according to the preceding lemma, we have
$|L(E)-L(E^{\prime})|\leqslant 3C_{\ell,I}\Gamma^{2n-1}|E-E^{\prime}|+C_{0}e^{-c_{0}n}$
for all $n\in{\mathbb{N}}$. Put differently, there is a constant $\tilde{C}>1$ such that
$|L(E)-L(E^{\prime})|\leqslant(\tilde{C})^{n}|E-E^{\prime}|+\tilde{C}e^{-c_{0}n}$
Choosing $n=\Bigl[\frac{\ln|E-E^{\prime}|^{-1}}{3\ln\tilde{C}}\Bigr$], we obtain
$|L(E)-L(E^{\prime})|\leqslant|E-E^{\prime}|^{2/3}+\tilde{C}|E-E^{\prime}|^{\frac{c_{0}}{3\ln\tilde{C}}}$.
The proof is complete. $\Box$